### v1

# Efficiently serving large multimedia models using EPD Disaggregation

Abstract
Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead. This step helps convert raw inputs into tokenized representations that inflate the token sequence for the prefill phase, negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput. We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources. Unlike current systems, which bundle encoding and prefill together, our disaggregation approach alleviates memory bottlenecks, mitigates synchronization delays, and supports flexible batching. Specifically, we employ a new caching mechanism for multimodal tokens, enabling asynchronous transfer of multimodal tokens and introduce an integrated module to find optimal config for EPD system and minimize resource usage while maximizing SLO-based performance metric. Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ lesser for encoding-stage GPUs), that supports upto 22$\times$ higher batch sizes, 10$\times$ more number of images/ request, 2.2$\times$ higher kv cache size. Further, it leads to significant improvements in end-to-end throughput (up to 57% better), and latency metrics (TTFT up to 71% lower), compared to systems that do not disaggregate. Our findings underscore the potential of EPD disaggregation to enable resource-efficient and high-performance multimodal inference at scale.
1 Introduction
Large Language Models (LLMs) have revolutionized language understanding and reasoning, achieving superhuman performance across a variety of tasks [1, 4]. Recently, the scope of these models has expanded to include multiple modalities, such as images, audio, and videos, resulting in the advent of Large Multimodal Models (LMMs) [21, 11, 5]. LMMs enable users to interact with diverse data types, such as posing questions about visual scenes or analyzing audio clips, thereby unlocking novel applications across fields like healthcare, autonomous systems, and creative industries.
However, serving LMMs in an efficient manner presents unique challenges. Meeting strict Service Level Objectives (SLOs) such as end-to-end throughput (E2ETP), time to first token (TTFT), and time per output token (TPOT) becomes increasingly difficult given the additional computational and memory demands introduced by processing multimodal data [12]. Unlike LLMs, where inference involves prefill and decoding stages, LMMs require an additional encoding stage to process raw multimodal inputs (e.g., images or videos) into tokenized representations. This stage is computationally intensive, especially for high-resolution or complex multimodal inputs, and often produces a substantial number of additional tokens [20]. The resulting increase in tokens inflates resource consumption and leads to quadratic growth in prefill-stage compute demands, adversely impacting SLO attainment.
Disaggregating the prefill stage from the decode stage has emerged as a well-studied solution in the literature to improve inference efficiency for LLMs [24, 15, 14, 9, 7]. By assigning separate resources to each stage, prefill-decode disaggregation enables independent optimization of batching, scheduling, and resource allocation strategies, significantly enhancing system throughput and memory utilization. However, these techniques fail to address the challenges of LMM deployment, where the addition of an encoding stage fundamentally alters the resource dynamics. The encoding stage, with its high computational and memory overhead, introduces token inflation and dependency bottlenecks that ripple across subsequent stages, necessitating a rethinking of disaggregation strategies to accommodate multimodal workloads.
In current systems, the encoding and prefill stages are typically aggregated into a single monolithic and synchronous step executed on the same set of GPUs. This approach assumes that combining encoding and prefill can simplify implementation, but in practice, it exacerbates inefficiencies such as memory bottlenecks, limited batching flexibility, and synchronization delays. Specifically, simultaneous loading of both the multimodal encoder and LLM onto the same GPUs restricts memory for higher-resolution inputs or larger batch sizes, while the sequential execution of encoding and prefill introduces delays for encoding-light requests that must wait for encoding-heavy ones in the same batch. As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, making it clear that current solutions are inadequate for LMM workloads.
To address these challenges, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encoding, prefill, and decode stages, assigning each with dedicated resources to operate independently. This disaggregation enables customized batching, parallelization, and scheduling strategies for each stage, optimizing resource utilization while reducing contention. By separating the multimodal encoder from the LLM on distinct GPUs, the system significantly improves memory efficiency, allowing for larger batch sizes and higher-resolution inputs. Additionally, asynchronous token transfer between stages minimizes latency, ensuring smooth and efficient processing. As a result, EPD disaggregation achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT, TPOT, and E2ETP.
The major contributions of this work are as follows:
  • We identify major challenges and performance bottlenecks due to which the performance of existing multimodal inference systems is suboptimal. Specifically, these stem from encoding and prefill being a single monolithic and synchronous step that results in poor performance owing to inefficient utilization of memory and computation resources.
  • We propose an efficient system for LMM inference that relies on the novel idea of disaggregating encoding and prefill stages. Our approach addresses key challenges of resource allocation, inter-stage communication, and performance optimization within this framework. Specifically, we propose allocation of a new cache to store multimodal tokens on encoding and prefill stage that allows asynchronous transfer of multimedia tokens in an efficient manner.
  • We formulate the search for the optimal configuration of the EPD disaggregated system as a black-box optimization problem. This approach is simple yet practical because, on the one hand, it allows for a flexible search space that easily incorporates realistic configurations, including various parallelization strategies, GPU allocation, batch size selection, and complex scheduling strategies. On the other hand, it is computationally efficient, as a simulation outputs performance metrics (e.g., goodput or E2ETP) without actually deploying the system.
  • We provide a detailed experimental analysis of our method with respect to baselines on various popular LMM models, namely MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B. The results demonstrate the superiority of our method in terms various aspects, like memory (upto 15x lesser), facilitating upto 22x higger batch sizes, 10x more number of images/ request, 2.2x higher kv cache size. Further, our method achieves significantly improved performance in terms of SLO attainment (upto 90-100% more), TTFT latency (upto 71% lesser), and E2ETP (upto 57% better).
2 Related Work
Serving LLMs and LMMs, which often consist of billions of parameters, poses significant challenges due to slow inference speeds [12, 25]. Efficiently serving these models to billions of users while adhering to stringent Quality of Service (QoS) criteria has therefore become an active area of research [6]. This has led to the development of numerous production-grade systems, such as vLLM [10], TorchServe [17], NVIDIA Triton [13], SGLang [23], and HuggingFace Text Generation Inference (TGI) [18], that integrate ideas emerging from recent advancements in this field.
Throughput-Optimized Inference Systems: One prominent direction of research focuses on improving throughput, a critical system-level metric, by optimizing GPU utilization. For instance, Orca [22] proposed continuous batching, a technique that dynamically increases batch size without interrupting ongoing decoding, thereby enhancing GPU efficiency. vLLM [10] introduced paged KV cache management, which reduces memory fragmentation and frees up GPU memory, enabling larger batch sizes and ultimately boosting throughput. Similarly, SARATHI [2] introduced chunked prefill, which splits prefill requests into smaller chunks and piggybacks decoding requests alongside them to improve GPU utilization. However, these systems are designed for LLMs and do not account for the resource-intensive multimodal encoding step unique to LMMs, which precedes the prefill phase.
Latency-Optimized Inference Systems: Another line of work focuses on improving user-level metrics, such as request latency, for individual queries. For example, FastServe [19] employs iteration-level preemptive scheduling to reduce latency caused by head-of-line blocking in throughput-optimized systems. To gain finer control over TTFT and TPOT, approaches like SplitWise [14], DistServe [24], and DéjàVu [16] disaggregate the prefill and decode stages, mitigating interference between them. However, these methods are designed for LLMs and overlook the resource-intensive encoding step present in LMMs. Since encoding is coupled with prefill, these systems suffer from poor TTFT latency, particularly when the resolution and quantity of multimodal data are large.
We propose disaggregation of encoding from prefill, enabling fine-grained control over batching, scheduling, and parallelization strategies, thereby achieving better latency and throughput in LMMs.
3 Method
Figure 1: The inference pipeline of EPD Disaggregation system.
Figure 2: System architecture of the proposed EPD Disaggregated Inference. It consists of three stages. In each stage, data parallelism (DP) means multiple instances serve in parallel. Each instance encompasses pipeline-parallel (PP) workers and tensor-parallel (TP) workers. For the encoding stage, since intra-request parallel (IRP) is better than TP, we use IRP instead of TP workers. MM stands for Multi-Media.
3.1 LMM System
At a high level, the LMM system transforms multimodal data and a text prompt into a coherent textual output. The system receives three types of input: (1) a text prompt $i_{p}$, which provides the context for the task (e.g., “What is the content of the image?”), (2) multimodal data $i_{m}$, which includes images, videos, or other sensory inputs (e.g., one or more images of a cat), and (3) sampling data $i_{s}$, which consists of parameters such as the desired number of output tokens and temperature settings. The system processes these inputs to generate a text output $o$, which is the model’s response (e.g., “The image contains a cat.”). For simplicity, we focus on the visual modality in this paper, but the framework is extensible to other types of multimodal data as well.
3.2 EPD Disaggregation
Disaggregating an LMM system involves breaking the inference process into distinct stages, each with specific inputs and outputs. As shown in Figure  1, the system follows a multi-stage inference pipeline consisting of three primary components: encoding, prefill, and decode. Additionally, the transitions between these stages are managed through EP-migration and PD-migration, which handle the transfer of data between encoding-prefill and prefill-decode stages, respectively. In the following, we describe each stage and the transformations involved.
Encoding (E): The multimodal input $i_{m}$ is processed by the multimodal encoder (MME) $E$, which converts the input data into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs. These tokens represent a high-dimensional embedding of the input and serve as the foundation for the next stage of the pipeline.
$v_{t}^{e}=E(i_{m})$ (1)
EP-migration: Once the encoding stage $E$ completes processing, the generated multimodal tokens $v_{t}^{e}$ must be transferred to the prefill stage $P$. This transfer is managed by the EP-migration function $\psi_{EP}$, which handles the migration of tokens from the encoding worker’s cache to the prefill worker’s cache.
$v_{t}^{p}=\psi_{EP}(v_{t}^{e})$ (2)
Here, $v_{t}^{p}$ represents the tokens in the prefill stage after migration.
Prefill (P): The multimodal tokens $v_{t}^{p}$ arrive at the Prefill stage $P$, along with the text prompt $i_{p}$. In this stage, the system initializes the Key-Value (KV) cache $kv_{1}^{p}$ and generates the first output token $o_{1}^{p}$. The KV cache stores context and intermediate states, which are necessary for the subsequent decoding stage.
$kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$ (3)
PD-migration: After the prefill stage generates the initial token $o_{1}^{p}$ and updates the KV cache to $kv_{1}^{p}$, the data must be transferred to the decode stage $D$. This transfer is managed by the PD-migration function $\psi_{PD}$, which ensures that the KV cache and the generated token are correctly handed off to the decode worker.
$kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$ (4)
Decode (D): The Decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the previously generated token $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$. This autoregressive process continues until the desired output sequence is produced.
$kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$ (5)
Each token generation step depends on the previous token and the context stored in the KV cache, ensuring the system generates contextually accurate and coherent outputs.
3.3 System Design and Optimization
Figure 2 illustrates the architecture of our EPD Disaggregated Inference system. Each stage of the pipeline—Encoding, Prefill, and Decoding—features fully independent instances capable of running the corresponding stage. These instances are designed to operate in DP mode, meaning multiple instances per stage can process different requests concurrently, ensuring scalability and efficiency.
Each instance comprises a scheduler, responsible for scheduling requests, block managers (responsible for managing cache(s)), and multiple workers. The workers operate in tensor-parallel (TP) and/or pipeline-parallel (PP) mode, where each worker holds only a subset of the model weights and the corresponding caches required for the stage.
In the Encoding Stage, instances/ workers only load the MME as the LLM is not required at this stage. Similarly, only the MM cache is necessary to be initialized. In the Prefill Stage, instances load the LLM to handle prefill operations, but the MME is no longer required. These instances utilize both the MM cache and the KV cache to efficiently manage multimedia and key-value data associated with the request. In the Decoding Stage, instances also load the LLM, focusing exclusively on decoding tasks. The MME is not required at this stage, and only the KV cache is utilized.
All cache transfers occur when an instance pulls a request from its respective queue, ensuring that data is moved only when necessary. Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on reducing latency (Section 3.3.2), managing resources effectively (Section 3.3.3), and ensuring smooth token transfers between stages (Section 3.3.1).
3.3.1 Asynchronous Token Transfer
To minimize latency during token transfers between stages, our system supports direct, asynchronous token transfers via high-bandwidth NVLink/ InfiniBand / NVSwitch channels. The transfer occurs asynchronously, allowing the system to continue processing new requests while the token transfer is in progress. Specifically, we maintain a multimodal cache (MM cache) on both encoding and prefill workers. When a request completes the encoding stage, its tokens are first stored in the encoding worker’s MM cache, freeing the encoding worker to serve subsequent requests immediately.
An independent asynchronous event loop continually monitors for completed encoding tasks. When a request’s tokens become available, this event loop initiates a direct transfer of the tokens to the prefill worker’s MM cache. Because the transfer is asynchronous, the system can process other tasks in parallel without stalling. Once the token transfer is confirmed, the corresponding MM cache entries at the encoding stage are cleared to reclaim memory for future requests.
To manage these cache regions efficiently, we introduce an MMBlockManager. For each incoming request, the MMBlockManager precomputes the number of cache blocks the request will require and allocates them accordingly. After the request transitions from the encoding to the prefill stage and the tokens are transferred, those blocks are reassigned or de-allocated, enabling flexible cache usage even under heavy workloads.
Figure 3: The illustration of Intra-Request Parallel.
3.3.2 Intra-Request Parallel (IRP)
Often, multimodal requests include multiple high-resolution images, each of which is partitioned into patches for high-resolution encoding. For modern LMMs, the encoder responsible for processing these patches can be large, making the encoding phase computation-bound. Hence, if a request is processed serially, the encoding phase can become a bottleneck, adding significant delay to the first token latency TTFT.
To address this, we introduce Intra-Request Parallel (IRP), which splits a single request’s image patches among multiple encoding workers operating in data-parallel mode. We are guided by the intuition that these patches are encoded independently and hence can be processed in a parallel and asynchronous manner. As depicted in Fig. 3, each encoding worker receives a distinct subset of patches, computes the corresponding token representations, and transfers them to the prefill stage independently and asynchronously. This means that patches assigned to one encoding worker can finish and begin transferring without waiting for patches processed by other workers.
Once all the patch-level tokens arrive at the prefill stage, they are aligned, projected, and merged to form the complete set of multimodal tokens for the request. The request then proceeds through the prefill stage as usual, with the corresponding LLM processing the aggregated multimodal and prompt tokens. By allowing concurrent encoding of patches across multiple workers, IRP effectively reduces end-to-end latency and improves the utilization of available encoding resources, especially for high-resolution inputs with large numbers of patches.
3.3.3 Optimized Resource Allocation
The system dynamically determines configurations such as batch sizes, scheduling strategies, and parallelization approaches for each stage of the pipeline. This ensures that critical performance metrics are kept near optimal, based on the workload samples collected during deployment. These workload samples are gathered in real time from a recent time window. If the optimal performance metrics vary significantly in subsequent time windows, the system may adjust the configuration of the EPD pipeline accordingly.
We leverage a black-box optimization algorithm to select configurations, which minimizes GPU usage and maximizes a critical performance metric. For online serving scenarios, we define goodput as the key performance metric. Goodput represents the maximum request rate at which 90% of requests meet the TTFT and TPOT requirements. In offline scenarios, where TTFT and TPOT constraints are less critical, we use E2ETP as the performance metric. By the nature of performance metrics, the optimized config will minimize pipeline inefficiencies (e.g., idle time) and ensure allocated resources are utilized efficiently.
The optimization problem is formulated as:
$\max_{(\mathbf{p},\mathbf{b},\mathbf{s})\in\mathcal{X}}f(\mathbf{p},\mathbf{b},\mathbf{s})-\beta cost(\mathbf{p})$ (6)
where $\mathcal{X}$ is the search space for system configurations, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$. These variables are vectors, with each element representing the configuration of an individual instance. Here, an instance can handle requests independently, encompassing the (sub-)workers for tensor parallelism and pipeline parallelism. Since instances within a stage process different requests in parallel, we term it data parallelism. Note that $\mathbf{p},\mathbf{b},\mathbf{s}$ can have variable lengths, as the number of instances is itself a configurable parameter. For $i$-th instance, we denote its stage as $\text{Stage}_{i}\in\{E,P,D\}$.
  • Parallelization: Let $\mathbf{p}$ denote the vector of parallel configurations for all instances. For the $i$-th instance, if it is a prefill or decoding instance, then its configuration $p_{i}$ includes: $p^{\text{TP}}_{i}$, the number of GPUs used for tensor parallelism; and $p^{\text{PP}}_{i}$, the number of GPUs used for pipeline parallelism. If it is an encoding instance, considering intra-request parallism (IRP) does not require communication, which is better than TP, we only use IRP. Therefore, we overload the symbol $p^{\text{TP}}_{i}=p^{\text{IRP}}_{i}$ to denote the number of GPUs used for IRP. If the cost per GPU is a constant $c$, then the total cost is: $cost(\mathbf{p})=c\sum_{p_{i}\in\mathbf{p}}(p^{\text{TP}}_{i}\times p^{\text{PP}}_{i})$.
  • Max Batch Size determines how many requests are processed simultaneously during the encoding, prefill, and decoding stages. Let $\mathbf{b}$ denote the batch size config for all instances. For $i$-th instance, $b_{i}$ is its max batch size. This config involves a trade-off between latency and throughput. Larger batch sizes improve throughput by enabling parallel processing but may increase latency if the stages become compute-bound.
  • Scheduling involves two main questions: First, to which workers or queues should the requests be assigned? Second, how should the order of requests be determined within a queue? We adapt strategies from the DistServe framework [24]. In the encoding stage, when a request arrives, it is assigned and pushed to an instance queue. In the prefill and decode stages, a global queue is used, and each engine pulls from this queue when available. Possible assignment strategies in the encoding stage include Round-Robin or Least-Loaded First for assigning requests. Once a request is assigned to a worker’s queue, we can apply ordering strategies like first-come-first-serve or shortest-job-first, or more complex strategies that prioritize requests based on their Service Level Objectives (SLOs). For simplicity, we constrain that all instances within the same stage share the same scheduling strategy.
In the objective of the problem 6, $\beta$ is a weight that controls the trade-off between performance and cost reduction. In a black-box setting, the performance function $f(\cdot)$ is unknown and must be evaluated through simulations to obtain the performance metric values. To solve the optimization problem, existing methods such as random search, evolutionary algorithm, Bayesian optimization [3], and parametric iterated local search [8] can be employed to find the optimal configuration. We adopt random search to find a near-optimal config for its simplicity.
4 Implementation
EPD is a fully capable distributed serving system for LMMs, comprising several key components: a load estimation module, a resource allocation module, a RESTful API frontend, and a multimodal-aware orchestration layer. The entire framework is implemented with a mix of Python and C++/ CUDA implementations , ensuring superior scalability and performance. To facilitate integration, we repurpose the distributed execution engine from vLLM, which supports numerous popular LLMs and LMMs, allowing easy adaptation of new models into our disaggregated framework with minimal effort.
The API interface adheres to OpenAI’s multimodal specifications, enabling users to specify parameters such as output length, temperature, and multimodal data inputs.
The scheduler is specifically designed for the disaggregated EPD framework, dynamically managing batch sizes and enabling asynchronous execution of the encoding, prefill, and decoding phases. The load estimation module ensures efficient GPU allocation across these phases, adapting to changing workload demands in real time.
Our repurposed distributed execution engine uses Ray actors to implement GPU workers, which manage multimodal and key-value caches, and coordinate the independent execution of the encoding, prefill, and decoding tasks. Furthermore, it supports 3D parallelism, incorporating Data Parallelism (DP), Tensor Parallelism (TP), and Pipeline Parallelism (PP) to maximize resource utilization and scalability.
The orchestration layer includes custom CUDA kernels optimized for parallelism in the encoding and prefill phases. These kernels enable efficient management of paged multimodal caches and ensure seamless asynchronous transfer of multimodal tokens between encoding and prefill GPUs. The orchestration layer oversees the execution of encoding, prefill, and decoding instances, handling tasks such as request distribution, KV cache transmission, and result aggregation. For efficient data movement, the system employs NCCL for inter-node GPU communication and asynchronous CudaMemcpy for intra-node transfers, ensuring smooth operations without disrupting GPU computations across the disaggregated EPD framework.
5 Experiments
In this section, we analyze and compare the performance of the proposed EPD disaggregation method against the baselines.
Baseline Methods: We compared our proposed method EPD against two popular baselines: DistServe [24] and vLLM [10]. The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on a separate GPUs. In contrast, the vLLM baseline adopts a monolithic architecture, where all three stages runs on the same set of GPUs. In addition, we also consider an ablation of our method EPD without the IRP feature, named as EPD-IRP.
Since DistServe was originally designed for LLMs, we extended its framework to support LMMs. This required extending the request handling infrastructure to process multimodal data, modifying the block manager to accommodate multimodal tokens, and changing the execution engine from SwiftTransformer to vLLM, which natively supports multiple LLMs. The change to vLLM also ensured consistency in execution engine across all baselines, eliminating potential confounding performance differences arising from variations due to the execution engine.
Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [21], InternVL 2.0 8B, and InternVL 2.0 26B [5]. These models are renowned for their advanced capabilities in processing and understanding multimodal data.
MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in a total of 8 billion parameters. It excel in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [21].
InternVL 2.0 8B features an InternViT-300M-448px vision encoder with 300 million parameters, paired with an internlm2_5-7b-chat language model comprising 7.7 billion parameters, totaling 8 billion parameters. Similarly, InternVL 2.0 26B combines an InternViT-6B-448px-V1-5 vision encoder with 6 billion parameters and an internlm2-chat-20b language model containing 20 billion parameters, resulting in 26 billion parameters. These configurations result in models with 8 billion and 26 billion parameters, respectively.
Datasets: To analyze its performance across diverse scenarios, we utilized custom synthetic workload for all experiments. This approach offered the flexibility to systematically vary parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings. This controlled experimental setup allowed us to assess the system’s robustness and adaptability under different workload configurations, ensuring a comprehensive evaluation of its efficiency and scalability. The length of input tokens in the prompt (excluding multimodal tokens) was fixed to 22. The length of output tokens were fixed to 10 in end-to-end experiments and 1 in the first token generation experiments.
MiniCPMv 2.6 8B InternVL 8B InternVL 26B
#I/R TTFT TPOT TTFT TPOT TTFT TPOT
2 1.40 0.04 1.20 0.05 3.50 0.07
4 2.60 0.04 2.40 0.06 7.05 0.08
6 3.90 0.06 3.55 0.09 11.00 0.95
8 5.10 0.06 5.00 0.18 15.00 0.15
Table 1: Thresholds for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The runtime performance is assessed using the TTFT and TPOT metrics for online settings, and E2ETP for offline settings. These metrics are defined as follows:
  • TTFT: The time elapsed from the creation of a request in the Request Generator until the first token is received by the Decoder. This metric captures the responsiveness of the system to incoming requests.
  • TPOT: The time elapsed from the arrival of the first token at the Decoder to the arrival of the last token at the Output Consumer, divided by the number of output tokens. This metric reflects the per-token latency for generating outputs.
  • E2ETP: The end-to-end throughput of the system, measured in terms of requests completed per second. This metric evaluates the system’s ability to handle high request volumes in offline settings.
  • SLO Attainment: The percentage of requests that meet predefined Service Level Objectives (SLOs), such as TTFT and TPOT thresholds. This metric assesses the system’s ability to deliver consistent performance under operational constraints.
  • Goodput: The rate of successfully completed requests that meet SLO criteria, measured in requests per second. This metric emphasizes the system’s effective processing capacity, excluding failed or non-compliant tasks.
For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory. Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes. We demonstrate the advantages of our proposed approach using these memory-related metrics, highlighting its ability to enhance system scalability and performance.
Implementation Details We conducted our experiments using a cluster of 8 NVIDIA A100 GPUs (82GB), which provided the computational resources necessary for evaluating the proposed system. Each server was equipped with 128 CPUs and 1TB of RAM. The CUDA version was 12.2. Flash attention-2 was used for the attention implementation. We use FP16 precision for all experiments.
To ensure a fair comparison, we standardized key performance-affecting settings of the inference engine across all baselines. Specifically, these include a block size of 16; a maximum of 2048 blocks per request; context tokens capped at 49,152, and decoding tokens at 81,920 per batch. The scheduling policy for all stages was set to First-Come-First-Served (FCFS). Further, to allow enough resources for memory-heavy multimdodal requests to execute, KV cache GPU utililization was set to 50%, and the maximum number of multimedia data of 32 was imposed per prompt. The size of multimodal cache was fixed to 3000 across all models, and the vLLM inference engine was run in eager mode. Finally, for the vLLM inference engine, we used the version 0.6.1.post1 that represents a stable version for multimodal inference.
In our online experiments, requests were sent to the inference engine using a Poisson arrival process with a fixed $\lambda$, representing the number of requests per second. Each trial was executed until 100 requests were completed, ensuring sufficient data for consistent performance analysis. Further, to optimize TTFT and TPOT in this latency-sensitive setting, we disabled batching by setting batch sizes to 1 for the encoding, prefill, and decoding stages.
(a) MiniCPMv-2.6                                     (b) InternVL2-8B                                     (C) InternVL2-26B
Figure 4: SLO attainment for end-to-end inference across various models and image counts per request. Column (a) presents results for the MiniCPMv 2.6 model with 2, 4, 6, and 8 images per request (from top to bottom). Column (b) and (c) show results for InternVL2-8B and InternVL2-26B, respectively. The image resolution is fixed at 4K. As observed, EPD significantly outperforms all baselines across all configurations. EPD-IRP means that IRP is disabled in EPD system.
5.1 SLO Attainment for End-to-End Generation
In this experiment, we evaluate the end-to-end SLO performance of the proposed method (EPD), an ablation of our method without IRP (EPD-IRP) and existing state-of-the-art baselines, DistServe and vLLM. The setup involves 8 NVIDIA A100 GPUs serving 100 multimodal requests, arriving via a Poisson process with a fixed rate ($\lambda$) representing the number of requests per second. We tested three different models—MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B—under varying image counts per request All images were fixed at a 4K resolution of 4032x3024 pixels. The SLO criteria for each configuration, shown in Tab. 1, establish latency upper bounds for both TTFT and TPOT, ensuring timely token generation and consistent decoding performance.
The results for MiniCPMv 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right) are presented in Fig. 4. Each row corresponds to a different image count per request, specifically 2, 4, 6, and 8, listed from top to bottom. The X-axis represents the per-GPU request rate (requests per second, req/s), while the Y-axis shows the SLO attainment percentage, defined as the proportion of requests meeting both TTFT and TPOT thresholds. The black dotted line indicates the 90% SLO attainment threshold, and the corresponding request rate represents the goodput for 90% SLO attainment.
EPD-IRP demonstrates significantly higher SLO attainment compared to other baselines. At lower request rates, EPD-IRP achieves over 90% SLO attainment in most scenarios. This performance advantage is attributed to EPD-IRP’s ability to parallelize the computationally intensive image encoding step across multiple GPUs, mitigating the delay due to computationally-heavy encoding step. The second-best performer, EPD, achieves an SLO attainment close to 50% in many cases. However, the prior work, DistServe and vLLM, perform the worst, often attaining SLO percentages below 10%. These baselines suffer from interference between encoding and prefill stages, resulting in consistently poor SLO attainment. On the other hand, EPD can mitigate this interference due to disaggregation of encoding and prefill, however it lacks the parallelization ability of EPD-IRP.
As the image count per request increases (from top to bottom rows), the SLO attainment for all baselines declines, particularly at higher request rates. This decline reflects the growing computational demands of processing multiple high-resolution images within a single request. In the most demanding scenario, with 16 images per request (bottom-most row), SLO attainment drops significantly for all baselines. However, EPD-IRP continues to maintain reasonable SLO attainment, especially at lower request rates, highlighting its robustness under heavy workloads.
Finally, as we move from left to right across the plots, the change in model leads to progressively worse performance for all baselines. The InternVL models are prefill-heavy, with prefill stages taking substantially longer due to the generation of a higher number of image tokens. As a result, TTFT suffers from additional queuing delays at the prefill stage, especially at higher request rates. This effect becomes more pronounced when transitioning from the 8B (middle) to the 26B (right) model. In contrast, the MiniCPMv (left) model is optimized to generate fewer image tokens, making it encoding-heavy rather than prefill-heavy. Consequently, it avoids the queuing delays observed in the InternVL models and exhibits better latency characteristics overall.
(a) MiniCPMv-2.6                                     (b) InternVL2-8B                                     (C) InternVL2-26B
Figure 5: Distribution of TTFT (y-axis) for various #Images/ Request (x-axis) for the MiniCPMv 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right) models. Requests arrive following a Poisson distribution with a fixed request rate ($\lambda$). Note that, when comparing TTFT, vLLM is equivalent to DistServe, hence it is skipped. EPD-IRP means that IRP is disabled in EPD system.
5.2 First Token Generation Latency
In this experiment, we analyze the TTFT performance of the proposed method (EPD), an ablation of our method without IRP (EPD-IRP) and an existing state-of-the-art baseline, DistServe. Unlike prior analyses that focused on end-to-end inference, this experiment isolates the prefill phase of the computation, as the generation of the first token represents a critical latency bottleneck due to large number of prefill tokens in multimodel workloads. Note that in this setup the DistServe baseline is equivalent to vLLM since the decoding phase is not involved.
Fig. 5 shows the results for the three models: MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B, shown in the left, middle, and right plots, respectively. Each subplot presents box plots of TTFT distributions, with the x-axis representing the number of images per request and the y-axis showing the TTFT. The results are generated using 100 requests issued at a request rate of 0.25 for the MiniCPMv 2.6 and 0.08 for the InternVL2-8B and InternVL2-26B models. We limit the number of images per request (x-axis) on the basis of the respective model’s capacity.
As shown in the figure, EPD-IRP significantly outperforms the two baselines, EPD and DistServe, across all three models. Specifically, the TTFT can be upto 71.9%, 32.8%, and 44.9% lower than the DistServe baseline for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively. This performance advantage stems from its intra-request parallelization, which enables the multimedia data within each request to be encoded in parallel, whereas the baselines process this data sequentially.
Notably, EPD also demonstrates a slight performance improvement over DistServe, highlighting the independent benefits of disaggregating the encoding and prefill stages, even without full parallelization. In this case, the TTFT can be upto 9.8%, 4.9%, and 1.0% lower for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
Further, we also observe that the margin of improvement becomes increasingly pronounced as the number of images per request grows. These results highlight the ability of EPD-IRP to efficiently handle the computational demands of multimodal workloads, especially as the number of input images and overall complexity increase.
5.3 Memory Savings through Stage Disaggregation
This section demonstrates the significant memory savings achieved by disaggregating the encoding and prefill stages in the EPD system. In existing systems, such as DistServe and vLLM, the encoding and prefill stages are co-located on the same GPUs, requiring each GPU to load both the MME and the LLM. This co-location results in higher memory consumption, as both models must be loaded simultaneously on the same GPUs.
In contrast, the proposed disaggregated system optimizes memory usage by assigning the encoding stage to GPUs that load only the MME, while the prefill stage is handled by separate GPUs that load only the LLM. This separation significantly reduces memory requirements for both stages.
The memory savings achieved through this disaggregation are substantial. Specifically, the encoding-stage GPUs experience memory reductions of approximately 95%, 96.2%, and 78.3% for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively. On the other hand, the prefill-stage GPUs achieve memory savings of about 5%, 3.7%, and 21.6% for the same models. In practice, the memory savings could be as high as 15x lower as encoding stage do not need to load the KV cache which is only required for prefill. These reductions in memory usage enable the EPD system to allocate resources more efficiently, allowing for the support of additional tasks and models, as demonstrated in the subsequent experiments.
EPD Supports a Higher Number of Images per Request: Tab. 2 compares the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B. The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and 80% of the available memory allocated for KV cache.
For the MiniCPMv model, the EPD approach demonstrates a significant advantage in handling a higher number of images per request. At a resolution of 313x234, EPD supports up to 490 images per request, compared to just 77 for the aggregated DistServe system, resulting in a more than 6x improvement. At the 787x444 resolution, EPD can handle 165 images per request, which is a 6.3x increase over DistServe’s limit of 26. Even at the highest resolution (4032x3024), EPD supports 49 images per request, while DistServe is limited to only 7, demonstrating a 7x improvement.
For the InternVL2-26B model, the improvements are equally striking. EPD supports up to 10x more images per request at both the 313x234 and 4032x3024 resolutions, and approximately 4x more at the 787x444 resolution, compared to DistServe.
Interestingly, for the InternVL2-8B model, both EPD and DistServe support the same number of images per request across all resolutions (19). This is because the InternVL2-8B model model uses a relatively simple MME that is not memory-intensive, and the limit of 19 images is constrained by the maximum context length of the LLM rather than memory usage of the MME.
Model Image Reso. DistServe EPD
MiniCPMv 2.6 313,234 77 490
787,444 26 165
4032,3024 7 49
InternVL2-8B 313,234 19 19
787,444 19 19
4032,3024 19 19
InternVL2-26B 313,234 1 10
787,444 11 45
4032,3024 1 10
Table 2: Comparison of maximum supported number of images/request. Batch size is fixed to 1. KV cache is allocated as 80% of the available free memory. Experiment done on A800 GPU with 82GB memory. Higher (italicized) is better.
EPD Supports Higher Batch Sizes for Encoding and Prefill:
Tab. 3 compares the maximum supported batch sizes for the encoding (E) and prefill (P) stages across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B, at three different image resolutions. The number of images per request is fixed at 10, and each model uses a different patching strategy, leading to varying numbers of image patches. The last three columns show the maximum batch size that can be supported for each configuration without encountering out-of-memory (OOM) errors. For the EPD approach, batch sizes for encoding and prefill are listed separately, as these processes run independently on separate GPUs. In contrast, for the DistServe baseline, encoding and prefill share the same batch size since both stages are executed sequentially on a single GPU.
The EPD approach significantly outperforms the DistServe baseline in terms of supported batch sizes. For instance, with the InternVL2-26B model at a resolution of 787x444, EPD supports a batch size of 22 for encoding, while the DistServe baseline can only handle a batch size of 1, representing a 22x improvement. On the other hand, with the MiniCPMv2.6 model at a resolution of 787x444, EPD supports a batch size of 29 vs 2 for the DistServe baseline, representing a 14.5x improvement for prefill.
DistServe EPD
Model Image Reso. #Patch E, P E P
MiniCPMv 2.6 313,234 1 7 49 86
787,444 3 2 16 29
4032,3024 10 OOM 4 9
InternVL2-8B 313,234 13 2 15 2
787,444 3 9 67 10
4032,3024 13 2 15 2
InternVL2-26B 313,234 13 OOM 6 1
787,444 3 1 22 4
4032, 3024 13 OOM 6 1
Table 3: Comparison of maximum supported E and P batch size. As seen, E, P batch sizes is much higher in EPD system in comparison to DistServe. Number of images/ request is fixed to 10. KV cache is allocated as 80% of the available free memory. Experiment done on one A800 GPU with 82GB memory. Higher (italicized) is better.
EPD Supports Larger KV Cache Sizes: Tab. 4 compares the maximum KV cache size that can be allocated in the disaggregated (EPD) system with that of the aggregated (DistServe or vLLM) system. In this experiment, the batch size is fixed at 1, and the image resolution is set to 4032×3024. We evaluate two LMM models, InternVL and MiniCPMv, with varying numbers of images per request.
The results clearly demonstrate that EPD supports significantly larger KV cache sizes than DistServe. For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for Distserve, representing a 2.2x improvement.
Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an Out of Memory (OOM) error, indicating that it cannot process 20 images even with a KV cache size of 0. Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
Model # Images/Req. DistServe EPD
MiniCPMv 2.6 5 86% 99%
10 74% 97%
20 49% 95%
40 OOM 92%
80 OOM OOCL
InternVL2-8B 5 94% 95%
10 89% 91 %
20 OOCL OOCL
InternVL2-26B 5 67% 89%
10 36% 80%
20 OOM 63%
40 OOM OOCL
Table 4: Comparison of maximum supported KV cache size (in terms of percentage of free memory) on prefill node for various #images/ request. Image resolution fixed to 4K. OOM refers to Out of Memory and OOCL refers to Out of Context Limit. Experiments are done on A800 GPU with 82GB memory. Higher (italicized) is better.
Figure 6: Left: The impact of tuning the number of encoding workers in the EPD method. The annotation ‘xEyP‘ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P‘ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests. Middle: The impact of the number of images per request in the workload. Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
5.4 Throughput in Offline Settings
In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the encoding (E) and prefill (P) stages.
Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput. In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory. In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory. In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory.
In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
We use 8 A800 GPUs for this experiment. For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively. In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively. The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens.
We tune the system’s hyperparameters as outlined in Sec. 3.3.3 and evaluate their impact on throughput. The results are presented in Fig. 6.
The left plot illustrates the importance of selecting an appropriate GPU configuration. As described in Sec. 3.3.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput.
The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck.
The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method. This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec. 3.3.3 to automatically select an optimal batch size.
6 Conclusion
In this paper, we present a novel approach to optimizing Large Multimodal Model (LMM) systems through the disaggregation of key processing stages. By separating the encoding, prefill, and decoding tasks into distinct stages, our system offers enhanced flexibility in resource allocation, enabling more efficient management of computational and memory resources. This disaggregation, combined with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage.
Our results demonstrate significant improvements in both throughput and memory efficiency, particularly in heterogeneous environments where resources are limited. By efficiently parallelizing the encoding and prefill stages, we are able to mitigate common bottlenecks and improve the scalability of multimodal inference tasks. Furthermore, the optimized memory management strategies ensure that the system can scale to handle larger models and more complex multimodal inputs, which are increasingly common in real-world applications.
The approach presented in this work represents a significant step forward in the development of scalable and efficient systems for multimodal AI, offering a foundation for future research and practical applications in areas such as natural language processing, computer vision, and cross-modal understanding. In summary, our system provides a robust solution for improving the performance of LMMs, making them more efficient and adaptable for deployment in real-world, resource-constrained environments.

### v2

@@
-Efficiently serving large multimedia models using EPD Disaggregation
+Efficiently Serving Large Multimodal Models Using EPD Disaggregation
 Abstract
 Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead.
-This step helps convert raw inputs into tokenized representations that inflate the token sequence for the prefill phase, negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput.
+This step negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput (E2ETP).
 We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources.
-Unlike current systems, which bundle encoding and prefill together, our disaggregation approach alleviates memory bottlenecks, mitigates synchronization delays, and supports flexible batching.
-Specifically, we employ a new caching mechanism for multimodal tokens, enabling asynchronous transfer of multimodal tokens and introduce an integrated module to find optimal config for EPD system and minimize resource usage while maximizing SLO-based performance metric.
-Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ lesser for encoding-stage GPUs), that supports upto 22$\times$ higher batch sizes, 10$\times$ more number of images/ request, 2.2$\times$ higher kv cache size.
-Further, it leads to significant improvements in end-to-end throughput (up to 57% better), and latency metrics (TTFT up to 71% lower), compared to systems that do not disaggregate.
-Our findings underscore the potential of EPD disaggregation to enable resource-efficient and high-performance multimodal inference at scale.
+Unlike current systems, which bundle encoding and prefill together, our approach decouple these steps unlocking new opportunities and optimizations.
+These include a new mechanism to cache multimedia tokens for efficient transfer, a novel way to parallelize encoding load within a request, a module to find the optimal resource allocation for disaggregated serving, and a novel role switching method to handle changing workload characteristics.
+Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ less utilization), batch sizes (up to 22$\times$ larger), 10$\times$ more images/request, and 2.2$\times$ larger KV caches.
+Further, it leads to significant improvements in latency metrics (TTFT up to 71% reduction) and end-to-end throughput (up to 57% reduction), compared to systems that do not disaggregate.
 1 Introduction
@@
 LMMs enable users to interact with diverse data types, such as posing questions about visual scenes or analyzing audio clips, thereby unlocking novel applications across fields like healthcare, autonomous systems, and creative industries.
+Figure 1: Above: Aggregated system, where E and LLM are coupled and must run on the same GPU.
+Below: Disaggregated system, where E runs on the 1st GPUs and LLM runs on the rest of three GPUs.
+Multimodal encoder and LLM are denoted by E and LLM, respectively.
 However, serving LMMs in an efficient manner presents unique challenges.
@@
 The encoding stage, with its high computational and memory overhead, introduces token inflation and dependency bottlenecks that ripple across subsequent stages, necessitating a rethinking of disaggregation strategies to accommodate multimodal workloads.
-In current systems, the encoding and prefill stages are typically aggregated into a single monolithic and synchronous step executed on the same set of GPUs.
-This approach assumes that combining encoding and prefill can simplify implementation, but in practice, it exacerbates inefficiencies such as memory bottlenecks, limited batching flexibility, and synchronization delays.
-Specifically, simultaneous loading of both the multimodal encoder and LLM onto the same GPUs restricts memory for higher-resolution inputs or larger batch sizes, while the sequential execution of encoding and prefill introduces delays for encoding-light requests that must wait for encoding-heavy ones in the same batch.
-As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, making it clear that current solutions are inadequate for LMM workloads.
-To address these challenges, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encoding, prefill, and decode stages, assigning each with dedicated resources to operate independently.
+Figure 2: Impact of disaggregation on batch size and number of images/ request for the MiniCPM-V 2.6 model
+These challenges of LLM deployment present opportunities for optimization that current serving systems do not consider.
+Currently, the encode and prefill stages are aggregated into a single monolithic and synchronous step executed on the same set of GPUs.
+Sequential execution of encoding and prefilling introduces interference.
+This interference is demonstrated graphically in Figure 1.
+In the aggregated approach (Figure 1 top), the prefill step ($LLM^{4}$) interferes with the encode-heavy request ($E^{5}$).
+As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, revealing that current solutions are inadequate for LMM workloads.
+By disaggregating encode from prefill, the system may be optimized to reduce interferences for specific workloads.
+From another perspective, simultaneous loading of both the Multimodal encoder (MME) and LLM onto the same GPUs in aggregated setups restricts memory for processing multiple high resolution images or larger batch sizes.
+In preliminary investigations, by removing the LLM and associated KV cache, we observe a significant increase in maximum batch sizes and supported number of images/ request (Figure 2).
+Thus, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encode, prefill, and decode stages, assigning each with dedicated resources to operate independently.
 This disaggregation enables customized batching, parallelization, and scheduling strategies for each stage, optimizing resource utilization while reducing contention.
-By separating the multimodal encoder from the LLM on distinct GPUs, the system significantly improves memory efficiency, allowing for larger batch sizes and higher-resolution inputs.
-Additionally, asynchronous token transfer between stages minimizes latency, ensuring smooth and efficient processing.
 As a result, EPD disaggregation achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT, TPOT, and E2ETP.
 The major contributions of this work are as follows:
-  • We identify major challenges and performance bottlenecks due to which the performance of existing multimodal inference systems is suboptimal.
-Specifically, these stem from encoding and prefill being a single monolithic and synchronous step that results in poor performance owing to inefficient utilization of memory and computation resources.
   • We propose an efficient system for LMM inference that relies on the novel idea of disaggregating encoding and prefill stages.
-Our approach addresses key challenges of resource allocation, inter-stage communication, and performance optimization within this framework.
-Specifically, we propose allocation of a new cache to store multimodal tokens on encoding and prefill stage that allows asynchronous transfer of multimedia tokens in an efficient manner.
-  • We formulate the search for the optimal configuration of the EPD disaggregated system as a black-box optimization problem.
-This approach is simple yet practical because, on the one hand, it allows for a flexible search space that easily incorporates realistic configurations, including various parallelization strategies, GPU allocation, batch size selection, and complex scheduling strategies.
-On the other hand, it is computationally efficient, as a simulation outputs performance metrics (e.g., goodput or E2ETP) without actually deploying the system.
-  • We provide a detailed experimental analysis of our method with respect to baselines on various popular LMM models, namely MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B.
-The results demonstrate the superiority of our method in terms various aspects, like memory (upto 15x lesser), facilitating upto 22x higger batch sizes, 10x more number of images/ request, 2.2x higher kv cache size.
-Further, our method achieves significantly improved performance in terms of SLO attainment (upto 90-100% more), TTFT latency (upto 71% lesser), and E2ETP (upto 57% better).
+Our approach addresses key challenges of inter-stage communication, efficient parallelization, resource allocation, and performance optimization within this framework.
+  • We propose intra-request parallelization (IPR) that shards a request into independent encoding jobs which can be executed in parallel, significantly reducing first token latency.
+  • We formulate the resource allocation problem as an optimization over batch sizes, scheduling strategies, and parallelization approaches for each pipeline stage.
+Further, we provide a simple black box optimizer that uses workload samples to approximate the optimal config for a given workload.
+  • We develop dynamic role switching capability that monitors the system for any bottlenecks and allows efficient reallocation of resources between E, P and D stages by efficiently switching ($\sim$0.7s) the role of an instance among these stages.
+This ensures system can respond to any changes in workload requirements during online serving.
+  • We conduct evaluations on various popular LMM models, namely MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B using real-world and synthetic workloads.
+The results demonstrate the superiority of our method across various aspects.
+For instance, our approach requires up to 15x less memory, facilitating up to 22x larger batch sizes, 10x more images/request, and 2.2x larger KV caches.
+Further, our method significantly improves performance in terms of SLO attainment (up to 90–100% more) and TTFT latency (up to 71% reduction).
 2 Related Work
-Serving LLMs and LMMs, which often consist of billions of parameters, poses significant challenges due to slow inference speeds [12, 25].
-Efficiently serving these models to billions of users while adhering to stringent Quality of Service (QoS) criteria has therefore become an active area of research [6].
-This has led to the development of numerous production-grade systems, such as vLLM [10], TorchServe [17], NVIDIA Triton [13], SGLang [23], and HuggingFace Text Generation Inference (TGI) [18], that integrate ideas emerging from recent advancements in this field.
-Throughput-Optimized Inference Systems: One prominent direction of research focuses on improving throughput, a critical system-level metric, by optimizing GPU utilization.
-For instance, Orca [22] proposed continuous batching, a technique that dynamically increases batch size without interrupting ongoing decoding, thereby enhancing GPU efficiency.
-vLLM [10] introduced paged KV cache management, which reduces memory fragmentation and frees up GPU memory, enabling larger batch sizes and ultimately boosting throughput.
-Similarly, SARATHI [2] introduced chunked prefill, which splits prefill requests into smaller chunks and piggybacks decoding requests alongside them to improve GPU utilization.
-However, these systems are designed for LLMs and do not account for the resource-intensive multimodal encoding step unique to LMMs, which precedes the prefill phase.
-Latency-Optimized Inference Systems: Another line of work focuses on improving user-level metrics, such as request latency, for individual queries.
-For example, FastServe [19] employs iteration-level preemptive scheduling to reduce latency caused by head-of-line blocking in throughput-optimized systems.
-To gain finer control over TTFT and TPOT, approaches like SplitWise [14], DistServe [24], and DéjàVu [16] disaggregate the prefill and decode stages, mitigating interference between them.
-However, these methods are designed for LLMs and overlook the resource-intensive encoding step present in LMMs.
-Since encoding is coupled with prefill, these systems suffer from poor TTFT latency, particularly when the resolution and quantity of multimodal data are large.
-We propose disaggregation of encoding from prefill, enabling fine-grained control over batching, scheduling, and parallelization strategies, thereby achieving better latency and throughput in LMMs.
+In this section, we describe feasible approaches in the literature for multi-modality model serving and their limitations and then review the use of the disaggregation technique.
+Multi-modality Model Serving.
+There are two ways to enable a multi-modality model serving system: (1) adopting and extending existing LLM serving systems, such as vLLM [9], SARATHI [2], and Orca [20]; and (2) leveraging the open-sourced code from recent works of improving multi-modality model inference.
+The first way fails to address the encoding bottleneck inherent in LMMs, limiting their applicability to multimodal workloads, particularly when handling high-volume and high-resolution multi-media data.
+In the second way, recent advancements such as KV cache eviction [10, 13] and compression [8] focus on specific features and limited scenarios.
+For instance, Inf-MLLM [13] enables efficient streaming inference for LMMs on a single GPU, targeting resource-constrained scenarios.
+While their implementations are feasible for serving LMMs, they fail to meet user SLOs in cloud serving scenarios.
+To the best of our knowledge, we are the first to propose an LMM serving system that leverages disaggregation and a series of integrated techniques to enable better resource allocation and improved SLOs.
+Disaggregated Serving.
+Disaggregated serving has emerged as a promising technique in large model serving.
+By decoupling the prefill and decode stages, systems like SplitWise [14], DistServe [21], and DéjàVu [16] mitigate interference between these phases, enabling finer control over Time to First Token (TTFT) and Time Per Output Token (TPOT).
+However, these systems primarily target LLMs and overlook the encoding step required for LMMs, which is tightly coupled with prefill.
+Recent work, such as Mooncake [15] and PD-Serve [7], has extended disaggregation to include KV cache management and other system-level optimizations.
+Despite these advancements, they remain limited in their applicability to LMMs.
+Disaggregating the encoding phase from prefill could unlock new opportunities for optimizing LMM serving.
+For instance, when handling requests with long videos, encoding the frames could be executed in parallel, reducing head-of-line blocking and improving overall system efficiency.
+This approach would enable more flexible batching and scheduling strategies, ultimately achieving better latency and throughput for multimodal workloads.
 3 Method
-Figure 1: The inference pipeline of EPD Disaggregation system.
-Figure 2: System architecture of the proposed EPD Disaggregated Inference.
-It consists of three stages.
-In each stage, data parallelism (DP) means multiple instances serve in parallel.
-Each instance encompasses pipeline-parallel (PP) workers and tensor-parallel (TP) workers.
-For the encoding stage, since intra-request parallel (IRP) is better than TP, we use IRP instead of TP workers.
-MM stands for Multi-Media.
-3.1 LMM System
-At a high level, the LMM system transforms multimodal data and a text prompt into a coherent textual output.
-The system receives three types of input: (1) a text prompt $i_{p}$, which provides the context for the task (e.g., “What is the content of the image?”), (2) multimodal data $i_{m}$, which includes images, videos, or other sensory inputs (e.g., one or more images of a cat), and (3) sampling data $i_{s}$, which consists of parameters such as the desired number of output tokens and temperature settings.
-The system processes these inputs to generate a text output $o$, which is the model’s response (e.g., “The image contains a cat.”).
-For simplicity, we focus on the visual modality in this paper, but the framework is extensible to other types of multimodal data as well.
+In this section, first, we describe the EPD disaggregation framework (Sec.
+3.1), followed by a detailed explanation of the system design and optimizations (Sec.
+3.2), that includes the asynchronous token transfer, intra-request parallel, optimized resource allocation, and dynamic role switching.
+Figure 3: The inference pipeline of EPD Disaggregation.
 3.1 EPD Disaggregation
-Disaggregating an LMM system involves breaking the inference process into distinct stages, each with specific inputs and outputs.
-As shown in Figure  1, the system follows a multi-stage inference pipeline consisting of three primary components: encoding, prefill, and decode.
-Additionally, the transitions between these stages are managed through EP-migration and PD-migration, which handle the transfer of data between encoding-prefill and prefill-decode stages, respectively.
-In the following, we describe each stage and the transformations involved.
-Encoding (E): The multimodal input $i_{m}$ is processed by the multimodal encoder (MME) $E$, which converts the input data into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs.
-These tokens represent a high-dimensional embedding of the input and serve as the foundation for the next stage of the pipeline.
-$v_{t}^{e}=E(i_{m})$ (1)
-EP-migration: Once the encoding stage $E$ completes processing, the generated multimodal tokens $v_{t}^{e}$ must be transferred to the prefill stage $P$.
+As shown in Figure 3, disaggregating an LMM system involves breaking the inference process into multiple stages: encoding, prefill, and decode.
+Transitions between these stages—EP-migration and PD-migration—handle the transfer of data from encoding to prefill and from prefill to decode, respectively.
+Formally, we denote the input text prompt as $i_{p}$, multimodal data as $i_{m}$, and the text output as $o$, and provide details for each step and transition as follows:
+Encoding (E): The multimodal input $i_{m}$ is processed by the MME $E$, converting it into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs.
+These tokens form a high-dimensional multimodal embedding for the next pipeline stage.
+This process can be denoted as $v_{t}^{e}=E(i_{m})$.
+EP-migration: Once the $E$ stage completes processing, the generated multimodal tokens $v_{t}^{e}$ are transferred to the prefill stage $P$.
 This transfer is managed by the EP-migration function $\psi_{EP}$, which handles the migration of tokens from the encoding worker’s cache to the prefill worker’s cache.
-$v_{t}^{p}=\psi_{EP}(v_{t}^{e})$ (2)
+This process is expressed as $v_{t}^{p}=\psi_{EP}(v_{t}^{e})$.
 Here, $v_{t}^{p}$ represents the tokens in the prefill stage after migration.
@@
 The KV cache stores context and intermediate states, which are necessary for the subsequent decoding stage.
-$kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$ (3)
+The process is denoted by $kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$.
 PD-migration: After the prefill stage generates the initial token $o_{1}^{p}$ and updates the KV cache to $kv_{1}^{p}$, the data must be transferred to the decode stage $D$.
 This transfer is managed by the PD-migration function $\psi_{PD}$, which ensures that the KV cache and the generated token are correctly handed off to the decode worker.
-$kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$ (4)
-Decode (D): The Decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the previously generated token $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$.
-This autoregressive process continues until the desired output sequence is produced.
-$kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$ (5)
-Each token generation step depends on the previous token and the context stored in the KV cache, ensuring the system generates contextually accurate and coherent outputs.
+This is denoted as $kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$.
+Decode (D): The decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the token generated in the previous iteration $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$.
+This autoregressive process continues until the desired output sequence is generated, as denoted by $kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$.
+Figure 4: System architecture of the proposed EPD Disaggregated Inference.
 3.2 System Design and Optimization
-Figure 2 illustrates the architecture of our EPD Disaggregated Inference system.
-Each stage of the pipeline—Encoding, Prefill, and Decoding—features fully independent instances capable of running the corresponding stage.
-These instances are designed to operate in DP mode, meaning multiple instances per stage can process different requests concurrently, ensuring scalability and efficiency.
+Figure 4 illustrates the architecture of the EPD Disaggregated inference system.
+Each pipeline stage (Encoding, Prefill, and Decoding) has independent instances that run the corresponding stage.
+These instances operate in data parallel (DP) mode, enabling concurrent processing of multiple requests per stage and ensuring scalability and efficiency.
 Each instance comprises a scheduler, responsible for scheduling requests, block managers (responsible for managing cache(s)), and multiple workers.
 The workers operate in tensor-parallel (TP) and/or pipeline-parallel (PP) mode, where each worker holds only a subset of the model weights and the corresponding caches required for the stage.
-In the Encoding Stage, instances/ workers only load the MME as the LLM is not required at this stage.
-Similarly, only the MM cache is necessary to be initialized.
-In the Prefill Stage, instances load the LLM to handle prefill operations, but the MME is no longer required.
-These instances utilize both the MM cache and the KV cache to efficiently manage multimedia and key-value data associated with the request.
-In the Decoding Stage, instances also load the LLM, focusing exclusively on decoding tasks.
-The MME is not required at this stage, and only the KV cache is utilized.
-All cache transfers occur when an instance pulls a request from its respective queue, ensuring that data is moved only when necessary.
-Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on reducing latency (Section 3.3.2), managing resources effectively (Section 3.3.3), and ensuring smooth token transfers between stages (Section 3.3.1).
+In the encoding stage, workers load only the encoder weights and initialize the Multimodal (MM) cache.
+In the prefill stage, the LLM weights are loaded, and both the MM and KV caches are required to efficiently manage all the data associated with the request.
+In the decoding stage, workers load the LLM weights for decoding tasks and use the KV cache.
+Cache transfers occur asynchronously when an instance pulls a request from its queue, ensuring the transfer occurs once the downstream instance is ready.
+Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on ensuring smooth token transfers between stages (Section 3.2.1), reducing latency (Section 3.2.2), managing resources effectively (Section 3.2.3), and adapting to changing workloads (Section 3.2.4).
+The ablation of these features are shown in Sec.
+4.4.
 3.2.1 Asynchronous Token Transfer
-To minimize latency during token transfers between stages, our system supports direct, asynchronous token transfers via high-bandwidth NVLink/ InfiniBand / NVSwitch channels.
-The transfer occurs asynchronously, allowing the system to continue processing new requests while the token transfer is in progress.
-Specifically, we maintain a multimodal cache (MM cache) on both encoding and prefill workers.
-When a request completes the encoding stage, its tokens are first stored in the encoding worker’s MM cache, freeing the encoding worker to serve subsequent requests immediately.
-An independent asynchronous event loop continually monitors for completed encoding tasks.
-When a request’s tokens become available, this event loop initiates a direct transfer of the tokens to the prefill worker’s MM cache.
-Because the transfer is asynchronous, the system can process other tasks in parallel without stalling.
-Once the token transfer is confirmed, the corresponding MM cache entries at the encoding stage are cleared to reclaim memory for future requests.
-To manage these cache regions efficiently, we introduce an MMBlockManager.
-For each incoming request, the MMBlockManager precomputes the number of cache blocks the request will require and allocates them accordingly.
-After the request transitions from the encoding to the prefill stage and the tokens are transferred, those blocks are reassigned or de-allocated, enabling flexible cache usage even under heavy workloads.
-Figure 3: The illustration of Intra-Request Parallel.
+Disaggregating the system introduces an additional step of transferring vision tokens from the encoding to the prefill stage.
+To minimize latency during token transfers between stages, our system employs direct, asynchronous transfers via high-bandwidth channels (NVLink, and InfiniBand).
+The asynchronous transfer allows the system to continue processing new requests without interruption.
+Both encoding and prefill workers maintain an MM cache to facilitate this process.
+When encoding is complete, tokens are stored in the encoding worker’s MM cache, allowing it to serve new requests immediately.
+An asynchronous event loop monitors completed encoding tasks and initiates direct token transfers to the prefill worker’s MM cache.
+Once the transfer is confirmed, the encoding cache entries are cleared to free memory.
+To manage these cache blocks effectively, we introduce the MMBlockManager, which pre-allocates cache blocks based on each request’s needs.
+After token transfer, the blocks are reassigned or de-allocated, ensuring flexible cache utilization even under heavy workloads.
 3.2.2 Intra-Request Parallel (IRP)
-Often, multimodal requests include multiple high-resolution images, each of which is partitioned into patches for high-resolution encoding.
-For modern LMMs, the encoder responsible for processing these patches can be large, making the encoding phase computation-bound.
-Hence, if a request is processed serially, the encoding phase can become a bottleneck, adding significant delay to the first token latency TTFT.
-To address this, we introduce Intra-Request Parallel (IRP), which splits a single request’s image patches among multiple encoding workers operating in data-parallel mode.
-We are guided by the intuition that these patches are encoded independently and hence can be processed in a parallel and asynchronous manner.
-As depicted in Fig.
-3, each encoding worker receives a distinct subset of patches, computes the corresponding token representations, and transfers them to the prefill stage independently and asynchronously.
-This means that patches assigned to one encoding worker can finish and begin transferring without waiting for patches processed by other workers.
-Once all the patch-level tokens arrive at the prefill stage, they are aligned, projected, and merged to form the complete set of multimodal tokens for the request.
-The request then proceeds through the prefill stage as usual, with the corresponding LLM processing the aggregated multimodal and prompt tokens.
-By allowing concurrent encoding of patches across multiple workers, IRP effectively reduces end-to-end latency and improves the utilization of available encoding resources, especially for high-resolution inputs with large numbers of patches.
+Multimodal requests often include multiple high-resolution images in many practical scenarions like autonomous driving and video question answering.
+In modern LMMs, these images are further converted into a large number of patches that needs to be processed through a computationally-heavy MME.
+This significantly increases the computation load, making the encoding process a major bottleneck.
+To address this, we introduce Intra-Request Parallelism (IRP), which partitions a single request’s image patches across multiple encoding workers in a data-parallel fashion.
+Since patches are encoded independently, they can be processed and transferred concurrently.
+Specifically, each encoding worker concurrently processes a subset of patches, computes their token representations, and asynchronously transfers them to the prefill stage.
+Once all patch-level tokens reach the prefill stage, they are aligned, projected, and merged to form the complete multimodal tokens.
 3.2.3 Optimized Resource Allocation
-The system dynamically determines configurations such as batch sizes, scheduling strategies, and parallelization approaches for each stage of the pipeline.
-This ensures that critical performance metrics are kept near optimal, based on the workload samples collected during deployment.
-These workload samples are gathered in real time from a recent time window.
-If the optimal performance metrics vary significantly in subsequent time windows, the system may adjust the configuration of the EPD pipeline accordingly.
-We leverage a black-box optimization algorithm to select configurations, which minimizes GPU usage and maximizes a critical performance metric.
-For online serving scenarios, we define goodput as the key performance metric.
-Goodput represents the maximum request rate at which 90% of requests meet the TTFT and TPOT requirements.
-In offline scenarios, where TTFT and TPOT constraints are less critical, we use E2ETP as the performance metric.
-By the nature of performance metrics, the optimized config will minimize pipeline inefficiencies (e.g., idle time) and ensure allocated resources are utilized efficiently.
-The optimization problem is formulated as:
+The EPD system requires tuning various configurations, including batch sizes, scheduling strategies, and parallelization methods for each pipeline stage.
+To determine the optimal setup, we collect historical workload samples and apply a black-box optimizer.
+Let $f(\cdot)$ represent the system’s performance metric (e.g., goodput, see Section 4).
+Since $f(\cdot)$ is treated as a black-box function with an unknown internal mechanism, we rely on a simulator—extended from DistServe [21]—to evaluate performance metrics efficiently.
+The objective is to maximize performance while minimizing GPU usage, which inherently reduces pipeline inefficiencies (e.g., idle time) and improves resource utilization.
+Formally, we solve
 $\max_{(\mathbf{p},\mathbf{b},\mathbf{s})\in\mathcal{X}}f(\mathbf{p},\mathbf{b},\mathbf{s})-\beta cost(\mathbf{p})$ (1)
-where $\mathcal{X}$ is the search space for system configurations, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$.
-These variables are vectors, with each element representing the configuration of an individual instance.
-Here, an instance can handle requests independently, encompassing the (sub-)workers for tensor parallelism and pipeline parallelism.
-Since instances within a stage process different requests in parallel, we term it data parallelism.
-Note that $\mathbf{p},\mathbf{b},\mathbf{s}$ can have variable lengths, as the number of instances is itself a configurable parameter.
-For $i$-th instance, we denote its stage as $\text{Stage}_{i}\in\{E,P,D\}$.
-  • Parallelization: Let $\mathbf{p}$ denote the vector of parallel configurations for all instances.
-For the $i$-th instance, if it is a prefill or decoding instance, then its configuration $p_{i}$ includes: $p^{\text{TP}}_{i}$, the number of GPUs used for tensor parallelism; and $p^{\text{PP}}_{i}$, the number of GPUs used for pipeline parallelism.
-If it is an encoding instance, considering intra-request parallism (IRP) does not require communication, which is better than TP, we only use IRP.
-Therefore, we overload the symbol $p^{\text{TP}}_{i}=p^{\text{IRP}}_{i}$ to denote the number of GPUs used for IRP.
+Here, $\mathcal{X}$ is the search space for system configs, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$ (See Appendix A for details).
+We use Bayesian optimization [3] to solve Problem 1.
+3.2.4 Dynamic Role Switching
+The configuration optimizer described in the previous section can determine the optimal settings for a given workload.
+However, in an online environment, workload characteristics can change dynamically, requiring adjustments to the configuration.
+Re-initializing the entire system from scratch in response to these changes can be both difficult and costly.
+For instance, there may be ongoing requests in the encoding, context, or decoding phases with precomputed KV and/or VE caches stored in memory.
+A naive re-initialization would not only involve booting time but also force these requests to restart from the beginning, potentially causing a cascading impact on SLO metrics for subsequent requests.
+To overcome these challenges, we introduce dynamic role switching, which enables any instance in the E, P, or D stages to switch roles to any other stage (E, P, or D) with minimal overhead.
+At a high level, dynamic role switching continuously monitors the system’s queuing statistics across all stages and reallocates workers to stages experiencing higher demand.
+When a decision is made to transfer an instance from a source stage $S$ to a destination stage $T$, the migration process occurs in three key steps:
+  • Offload: The instance in the $S$ stage stops accepting new requests and redistributes its queued tasks to sibling instances in the same stage.
+  • Migration: The instance is reconfigured to meet the requirements of the $T$ stage.
+This may involve switching both the model and cache type.
+For example, if the E stage is involved, the instance may switch from an LLM to an MME model and from a KV cache to an MM cache.
+  • Onload: The migrated instance resumes processing new requests, helping to alleviate the queuing bottleneck at the $T$ stage.
+This process typically takes less than 0.7 seconds.
+The duration is longer for migrations involving the E stage due to model and cache changes, but significantly shorter when switching between P and D stages, as both the LLM and KV cache can be reused.
+4 Experiments
+In this section, we analyze and compare the performance of the proposed EPD disaggregation method against various baselines.
+We start with an end-to-end generation performance analysis in Sec.
+4.1, followed by an examination of first token latency in Sec.
+4.2.
+We then evaluate the memory savings achieved by EPD in Sec.
+4.3.
+Finally, an ablation analysis of key system components is presented in Sec.
+4.4.
+Additional experiments are detailed in the Appendix, including 1) a throughput comparison in offline settings (Appendix C); 2) an extension of our framework to Neural Processing Units (NPUs) demonstrating a 10% performance improvement (LABEL:sec:appendix-npu), and 3) additional SLO and memory experiments (Appendix D).
+(a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
+Figure 5: SLO attainment for end-to-end inference across various models and image counts per request.
+Columns (a), (b), and (c) correspond to the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively.
+The top and bottom rows are with 2 and 4 images per request, respectively.
+As observed, EPD significantly outperforms all baselines across all configurations.
+Baselines: We compared the proposed EPD method against two popular baselines: DistServe [21] and vLLM [9].
+The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on separate GPUs.
+Since DistServe was originally designed for LLMs, we extended it to support LMMs by enabling multimodal data processing and modifying the block manager to accommodate multimodal tokens.
+The vLLM baseline adopts a monolithic architecture, where all three stages run on the same set of GPUs.
+Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [19], InternVL2-8B, and InternVL2-26B [5].
+These LMMs are renowned for their advanced capabilities in processing and understanding multimodal data.
+Detailed descriptions of the LMMs and their sizes can be found in Appendix B.2.
+Datasets: To evaluate performance across diverse scenarios, we conducted experiments using two datasets.
+The first was a synthetic workload, that allowed setting custom parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings.
+Unless otherwise specified in our experiments, the input prompt length was set to 22 tokens.
+The second dataset was NextQA [18], a benchmark video question-answering dataset.
+Unlike synthetic workload, NextQA features human-annotated questions and answers, providing a more realistic reflection of real-world video request distributions.
+Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The related performance metrics are listed below.
+  • TTFT: The time from the submission of a request to the system until the first token is received by the user.
+  • TPOT: The average time interval between each output token (excluding the first token).
+  • SLO Attainment: The percentage of requests that meet predefined SLOs, such as TTFT and TPOT requirements.
+  • Goodput: The hightest request rate at which 90% or more SLO attainment is achieved.
+For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory.
+Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes.
+(See section 4.3)
+4.1 SLO Attainment for End-to-End Generation
+In this experiment, we evaluate the goodput of EPD and the baselines in an online setting, where 100 multimodal requests arrive following a Poisson process with rate $\lambda$.
+The length of output tokens is fixed to 10.
+We test requests with 2 and 4 images, each at a resolution of $4032\times 3024$.
+The details of the corresponding SLO criteria are discussed in Sec.
+B.3.
+The results for three models (columns) are presented in Fig.
+5.
+The X-axis represents the per-GPU request rate ($\lambda$/#GPUs), and the Y-axis indicates the percentage of requests meeting both TTFT and TPOT requirements, with a 90% threshold denoted by a black dotted line.
+(a) MiniCPMv-2.6                  (b) InternVL2-8B                     (c) InternVL2-26B
+Figure 6: Distribution of TTFT (Y-axis) for various #Images/ Request (X-axis) for the MiniCPM-V 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right)
+EPD outperforms all baselines, achieving over 90% SLO attainment at lower request rates.
+This is due to its ability to parallelize the computationally intensive image encoding step across multiple GPUs.
+DistServe and vLLM often maintain less than 10% attainment due to interference.
+Meanwhile, comparing the first and second rows for varying image counts per request (2 and 4), we observe that while more images per request increase workload, EPD maintains reasonable performance, whereas baseline SLO attainment drops significantly.
+Lastly, InternVL, being prefill-heavy, experiences queuing delays due to the higher number of image tokens, particularly at higher request rates.
+This effect exacerbates performance degradation from the 8B to the 26B model.
+In contrast, MiniCPM-V, optimized to generate fewer image tokens, avoids these delays and achieves better overall latency.
+Further, we provide more results with 6 and 8 images in Sec.
+D.1.
+Next, we repeat the experiment using the non-synthetic video question-answering dataset, NextQA [18].
+For this, we randomly sampled 100 examples, with input text token lengths ranging from 4 to 21 (average: 11.42) and output token lengths ranging from 1 to 7 (average: 2.75).
+Each video request was represented by 8 uniformly sampled frames, and we used MiniCPM-V 2.6 in the experiment, adhering to the SLO criteria of TTFT = 5.60 and TPOT = 0.06.
+As illustrated in Fig.
+7, EPD is the only framework achieving 90% SLO attainment at low request rates, demonstrating its superior effectiveness in handling diverse, real-world-like workloads compared to DistServe and vLLM.
+Figure 7: SLO attainment for end-to-end inference on the NextQA dataset. As seen, EPD outperforms other baselines by a significant margin.
+4.2 First Token Generation Latency
+Multimodal requests often place a heavy load on both the E and P phases, making the first token generation a key latency bottleneck.
+Therefore, in this experiment, we analyze the first token latency for various baselines.
+The results presented in Fig.
+6 shows box plots of TTFT distributions for three different models.
+Note that, as the decoding phase is excluded, the vLLM baseline is equivalent to DistServe and is thus skipped.
+Requests are generated according to a Poisson distribution with a fixed request rate $\lambda$.
+Specifically, $\lambda=0.25$ for the MiniCPM-V 2.6, and $\lambda=0.08$ for both the InternVL2-8B and InternVL2-26B.
+Thanks to intra-request parallelization, EPD significantly outperforms both baselines, EPD and DistServe.
+Specifically, TTFT is reduced by up to 71.9%, 32.8%, and 44.9% compared to the DistServe baseline for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
+The performance gain grows with the number of images per request, demonstrating EPD’s efficiency in handling heavy multimodal workloads.
+4.3 Memory Savings through Stage Disaggregation
+In this section we analyze the memory savings achieved by disaggregating the encoding and prefill stages.
+The E workers can save memory as they do not require LLM weights or the KV cache.
+Analyzing only the weight size indicates memory reduction of approximately 95%, 96.2%, and 78.3% for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models.
+Similarly, for the P workers, memory saving of about 5%, 3.7%, and 21.6% are achieved, respectively.
+In practice, since KV cache is also not required at E workers, the memory saving can be even higher (93.3% saving, i.e., 15$\times$ lower, according to profiling).
+These reductions in memory usage enable the EPD system to support higher numbers of images per request and batch sizes (discussed in the following experiments) and larger KV cache sizes (discussed in Appendix D.2).
+Model Image Reso. DistServe EPD
+MiniCPM-V 2.6 313,234 77 490
+787,444 26 165
+4032,3024 7 49
+InternVL2-8B 313,234 19 19
+787,444 19 19
+4032,3024 19 19
+InternVL2-26B 313,234 1 10
+787,444 11 45
+4032,3024 1 10
+Table 1: Comparison of maximum supported number of images/ request. Batch size is fixed to 1. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
+EPD Supports a Higher Number of Images per Request: We compare the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B.
+The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and with 80% of the available memory allocated for KV cache.
+As shown in Tab.
+1, EPD handles more images per request than DistServe, for example, at 4032$\times$3024 resolution, 7$\times$ more for InternVL2-26B and 10$\times$ more for InternVL2-8B.
+For InternVL2-8B, the limit of 19 images per request is due to its maximum context length.
+Without this constraint, a larger number of images per request could be supported.
+EPD Supports Higher Batch Sizes: We compare maximum supported batch sizes for E and P stages across different settings of image resolution and models in Tab.
+2.
+As seen, EPD significantly outperforms DistServe for both E and P batch sizes.
+For example, with InternVL2-26B at 787$\times$444 resolution, EPD supports a batch size of 22 for encoding vs.
+DistServe’s 1, achieving 22$\times$ improvement.
+Similarly, for MiniCPM-V 2.6 at 787$\times$444, EPD achieves a batch size of 29 vs.
+DistServe’s 2 for prefill, a 14.5$\times$ improvement.
+DistServe EPD
+Model Image Reso. #Patch (E, P) E P
+MiniCPMv 2.6 313,234 1 7 49 86
+787,444 3 2 16 29
+4032,3024 10 OOM 4 9
+InternVL2-8B 313,234 13 2 15 2
+787,444 3 9 67 10
+4032,3024 13 2 15 2
+InternVL2-26B 313,234 13 OOM 6 1
+787,444 3 1 22 4
+4032, 3024 13 OOM 6 1
+Table 2: Comparison of maximum supported E and P batch size. #images/ request is fixed to 10. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
+4.4 Ablation Study
+In this section, we analyze the impact of various components of the proposed system.
+#I/R=2 #I/R=4 #I/R=6 #I/R=8
+EPD 0.92 1.02 1.14 1.74
+w/o IRP 1.46 (1.6x) 2.47 (2.4x) 3.37 (2.9x) 4.27 (2.5x)
+Table 3: Effect of ablating IRP feature from the proposed system on TTFT (s).
+Disabling IRP negatively affects the TTFT (up to 2.9x worse) for various multiple images/ request (#I/R).
+Results are averaged over 100 requests.
+Effect of IRP: We analyze the effect of ablating IRP, presented in Section 3.2.2.
+The experimental settings are same as TTFT experiment in Sec.
+4.2.
+As shown in Tab.
+3, removing the IRP feature negatively affects TTFT across various images/ request.
+Moreover, the degradation exacerbates as the number of images/ request increases.
+This is because the IRP feature allows parallelization of encoding load within the same request across multiple GPUs.
+Goodput (r/s) $\uparrow$ TTFT (s) $\downarrow$ TPOT (s) $\downarrow$
+EDP 1.25 2.12 0.031
+w/o Opt. 0.56 (2.2x) 4.48 (2.1x) 0.025 (0.8x)
+Table 4: Effect of Offline Optimizer.
+When the configuration is set randomly, the goodput is 2.2× worse on average.
+$\downarrow$ indicates that lower is better and $\uparrow$ means that larger is better.
+Effect of Offline Optimizer: By default, the EPD system collects workload samples and finds the optimal config offline using the optimizer.
+To demonstrate its effect, we conduct an experiment without using the optimizer and set configs randomly.
+Specifically, we uniformly sample 10 configs at random and report the estimated expectation of the performance metric in the second row of Tab.
+4.
+For a fair comparison, when evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is the goodput of the EPD system.
+(See Appendix B.4 for more details) As shown in Tab.
+4, disabling the optimizer causes significant performance degradation, in terms of both goodput and TTFT.
+This demonstrates the necessity of the optimizer, especially when the EPD system has many tunable configs.
+Effect of Dynamic Role Switching: To analyze the impact of dynamic role switching, we conduct a controlled experiment simulating a shift in workload characteristics.
+Specifically, we generate 100 requests following the same static workload configuration as described in Section 3.2.4.
+However, to introduce an artificial workload change, we set the first 10 requests to generate 50 output tokens, while the remaining 90 requests generate 500 output tokens.
+The request arrival rate is fixed at 3 requests per second.
+The results are shown in Tab.
+5.
+Latency (s) $\downarrow$ TTFT (s) $\downarrow$ TPOT (s) $\downarrow$
+EPD 28.01 1.42 0.05
+w/o Switch 61.10 (2.2x) 1.33 (0.9x) 0.12 (2.4x)
+Table 5: Effect of ablating dynamic role switching feature from EPD.
+Disabling the worker migration feature negatively affects the TPOT (2.4x worse), which results in significantly poor end-to-end latency (2.2x worse).
+Results are averaged over 100 requests with one 4K image per request.
+$\downarrow$ indicates that lower is better.
+Without dynamic worker migration, the system performs poorly because it remains locked in the initial configuration (5E1P2D, optimized offline for 50 tokens) and is unable to adapt to the increased decoding demand.
+In contrast, the EPD system with migration dynamically reconfigures itself (2E1P5D) to handle the new workload (500 output tokens) by shifting 3 E instances to D, resulting in approximately 2× better performance.
+4.5 Adaptation to Neural Processing Units (NPUs)
+This section presents the results of adapting the proposed EPD framework to Huawei Ascend NPUs. For more details on the implementation and experiments, see Sec. B.5.
+First, we compare the EPD-NPU against the vLLM and DistServe baselines in terms of SLO attainment using the InternVL2-8B LMM.
+We used the same settings as mentioned in Sec.
+4.1, except that we used a heavy encoding workload of 8 4032x3024 images per request.
+The optimal configuration for this workload was found to be 5E2P1D and the corresponding results are presented in Figure 8.
+As seen, EPD is the only setting that can achieve the SLO requirements, while the other baselines are completely decimated even for low request rates.
+Figure 8: SLO curve for InternVL2-8B with the following requirements (TTFT=8.5 and TPOT=0.12) on the controlled experimental setup with 8 images per request.
+Next, we compare the TTFT of EPD-NPU against vLLM.
+We found that EPD-NPU demonstrated a 35.2% improvement over vLLM, which is significantly more than the 24.4% improvement for the GPU scenario.
+We attribute this additional $\sim$10to the proportionally heavier encode of LLMs on NPU compared to GPU.
+Please refer to Sec.
+D.3 for more details.
+This results in greater benefits of disaggregating the encode from prefill on NPU in comparison to GPU.
+5 Conclusion
+In this paper, we present a novel approach to optimizing Large Multimodal Model (LMM) systems through the disaggregation of key processing stages.
+By separating the encoding, prefill, and decoding tasks into distinct stages, our system offers enhanced flexibility in resource allocation, enabling more efficient management of computational and memory resources.
+This disaggregation, combined with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage.
+Appendix A Optimizer Details
+We recapitulate the configuration optimization problem:
+$\max_{(\mathbf{p},\mathbf{b},\mathbf{s})\in\mathcal{X}}f(\mathbf{p},\mathbf{b},\mathbf{s})-\beta\cdot\text{cost}(\mathbf{p})$ (2)
+For the search space $\mathcal{X}$, there could be implicit constraints during the configuration search.
+For example, the total number of GPUs must not exceed the available resources (e.g., 8 or 16).
+Alternatively, if the cloud service provider intends to fully utilize all available 8 GPUs, an implicit constraint could enforce the number of GPUs to be exactly 8.
+These constraints serve to reduce the search space.
+The system configuration involves parallelization configurations $\mathbf{p}$, maximum batch size configurations $\mathbf{b}$, and scheduling configurations $\mathbf{s}$.
+These variables are vectors, where each element corresponds to the configuration of an individual instance.
+Each instance is capable of handling requests independently, including managing (sub-)workers for tensor parallelism and pipeline parallelism.
+Instances within a stage process different requests in parallel, a concept referred to as data parallelism.
+Note that $\mathbf{p}$, $\mathbf{b}$, and $\mathbf{s}$ can have variable lengths, as the number of instances is itself a configurable parameter.
+For the $i$-th instance, we denote its stage as $\text{Stage}_{i}\in\{E,P,D\}$, where $E$ represents the encoder stage, $P$ represents the processing stage, and $D$ represents the decoder stage.
+  • Parallelization: Let $\mathbf{p}$ denote the vector of parallel configs for all instances.
+For the $i$-th instance, if it is a prefill or decoding instance, then its config $p_{i}$ includes: $p^{\text{TP}}_{i}$, the number of GPUs used for tensor parallelism; and $p^{\text{PP}}_{i}$, the number of GPUs used for pipeline parallelism.
+If it is an encoding instance, considering IRP does not require communication, which is better than TP, we only use IRP.
+Therefore for encode, we overload the symbol $p^{\text{TP}}_{i}=p^{\text{IRP}}_{i}$ to denote the number of GPUs used for IRP.
 If the cost per GPU is a constant $c$, then the total cost is: $cost(\mathbf{p})=c\sum_{p_{i}\in\mathbf{p}}(p^{\text{TP}}_{i}\times p^{\text{PP}}_{i})$.
@@
 Larger batch sizes improve throughput by enabling parallel processing but may increase latency if the stages become compute-bound.
-  • Scheduling involves two main questions: First, to which workers or queues should the requests be assigned?
-Second, how should the order of requests be determined within a queue?
-We adapt strategies from the DistServe framework [24].
+  • Scheduling involves two main decisions: First, which workers should each request be assigned to?
+Second, how should the order of requests be determined within a worker queue?
+To solve these decisions, we adapt strategies from the DistServe framework [21].
 In the encoding stage, when a request arrives, it is assigned and pushed to an instance queue.
-In the prefill and decode stages, a global queue is used, and each engine pulls from this queue when available.
+Between different stages, global queues are used, and each available engine pulls proactively from the queues.
 Possible assignment strategies in the encoding stage include Round-Robin or Least-Loaded First for assigning requests.
@@
 For simplicity, we constrain that all instances within the same stage share the same scheduling strategy.
-In the objective of the problem 6, $\beta$ is a weight that controls the trade-off between performance and cost reduction.
-In a black-box setting, the performance function $f(\cdot)$ is unknown and must be evaluated through simulations to obtain the performance metric values.
-To solve the optimization problem, existing methods such as random search, evolutionary algorithm, Bayesian optimization [3], and parametric iterated local search [8] can be employed to find the optimal configuration.
-We adopt random search to find a near-optimal config for its simplicity.
-4 Implementation
+Appendix B Implementation Details
 EPD is a fully capable distributed serving system for LMMs, comprising several key components: a load estimation module, a resource allocation module, a RESTful API frontend, and a multimodal-aware orchestration layer.
@@
 For efficient data movement, the system employs NCCL for inter-node GPU communication and asynchronous CudaMemcpy for intra-node transfers, ensuring smooth operations without disrupting GPU computations across the disaggregated EPD framework.
-5 Experiments
-In this section, we analyze and compare the performance of the proposed EPD disaggregation method against the baselines.
-Baseline Methods: We compared our proposed method EPD against two popular baselines: DistServe [24] and vLLM [10].
-The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on a separate GPUs.
-In contrast, the vLLM baseline adopts a monolithic architecture, where all three stages runs on the same set of GPUs.
-In addition, we also consider an ablation of our method EPD without the IRP feature, named as EPD-IRP.
-Since DistServe was originally designed for LLMs, we extended its framework to support LMMs.
-This required extending the request handling infrastructure to process multimodal data, modifying the block manager to accommodate multimodal tokens, and changing the execution engine from SwiftTransformer to vLLM, which natively supports multiple LLMs.
-The change to vLLM also ensured consistency in execution engine across all baselines, eliminating potential confounding performance differences arising from variations due to the execution engine.
-Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [21], InternVL 2.0 8B, and InternVL 2.0 26B [5].
-These models are renowned for their advanced capabilities in processing and understanding multimodal data.
-MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in a total of 8 billion parameters.
-It excel in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [21].
-InternVL 2.0 8B features an InternViT-300M-448px vision encoder with 300 million parameters, paired with an internlm2_5-7b-chat language model comprising 7.7 billion parameters, totaling 8 billion parameters.
-Similarly, InternVL 2.0 26B combines an InternViT-6B-448px-V1-5 vision encoder with 6 billion parameters and an internlm2-chat-20b language model containing 20 billion parameters, resulting in 26 billion parameters.
-These configurations result in models with 8 billion and 26 billion parameters, respectively.
-Datasets: To analyze its performance across diverse scenarios, we utilized custom synthetic workload for all experiments.
-This approach offered the flexibility to systematically vary parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings.
-This controlled experimental setup allowed us to assess the system’s robustness and adaptability under different workload configurations, ensuring a comprehensive evaluation of its efficiency and scalability.
-The length of input tokens in the prompt (excluding multimodal tokens) was fixed to 22.
-The length of output tokens were fixed to 10 in end-to-end experiments and 1 in the first token generation experiments.
-MiniCPMv 2.6 8B InternVL 8B InternVL 26B
-#I/R TTFT TPOT TTFT TPOT TTFT TPOT
-2 1.40 0.04 1.20 0.05 3.50 0.07
-4 2.60 0.04 2.40 0.06 7.05 0.08
-6 3.90 0.06 3.55 0.09 11.00 0.95
-8 5.10 0.06 5.00 0.18 15.00 0.15
-Table 1: Thresholds for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
-Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption.
-The runtime performance is assessed using the TTFT and TPOT metrics for online settings, and E2ETP for offline settings.
-These metrics are defined as follows:
-  • TTFT: The time elapsed from the creation of a request in the Request Generator until the first token is received by the Decoder.
-This metric captures the responsiveness of the system to incoming requests.
-  • TPOT: The time elapsed from the arrival of the first token at the Decoder to the arrival of the last token at the Output Consumer, divided by the number of output tokens.
-This metric reflects the per-token latency for generating outputs.
-  • E2ETP: The end-to-end throughput of the system, measured in terms of requests completed per second. This metric evaluates the system’s ability to handle high request volumes in offline settings.
-  • SLO Attainment: The percentage of requests that meet predefined Service Level Objectives (SLOs), such as TTFT and TPOT thresholds.
-This metric assesses the system’s ability to deliver consistent performance under operational constraints.
-  • Goodput: The rate of successfully completed requests that meet SLO criteria, measured in requests per second.
-This metric emphasizes the system’s effective processing capacity, excluding failed or non-compliant tasks.
-For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory.
-Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes.
-We demonstrate the advantages of our proposed approach using these memory-related metrics, highlighting its ability to enhance system scalability and performance.
-Implementation Details We conducted our experiments using a cluster of 8 NVIDIA A100 GPUs (82GB), which provided the computational resources necessary for evaluating the proposed system.
+B.1 Hyper-parameters in EPD system
+We conducted our experiments using a cluster of 8 NVIDIA A100 GPUs (82GB).
 Each server was equipped with 128 CPUs and 1TB of RAM.
@@
 The scheduling policy for all stages was set to First-Come-First-Served (FCFS).
-Further, to allow enough resources for memory-heavy multimdodal requests to execute, KV cache GPU utililization was set to 50%, and the maximum number of multimedia data of 32 was imposed per prompt.
-The size of multimodal cache was fixed to 3000 across all models, and the vLLM inference engine was run in eager mode.
-Finally, for the vLLM inference engine, we used the version 0.6.1.post1 that represents a stable version for multimodal inference.
+Further, to allow enough resources for memory-heavy multimodal requests to execute, KV cache GPU utilization was set to 50%, and the maximum number of multimedia data of 32 was imposed per prompt.
+The size of the multimodal cache was fixed to 3000 across all models, and the vLLM inference engine was run in eager mode.
+Finally, for the vLLM inference engine, we used the version 0.6.1.post1 which represents a stable version for multimodal inference.
 In our online experiments, requests were sent to the inference engine using a Poisson arrival process with a fixed $\lambda$, representing the number of requests per second.
@@
 Further, to optimize TTFT and TPOT in this latency-sensitive setting, we disabled batching by setting batch sizes to 1 for the encoding, prefill, and decoding stages.
+B.2 LMM Details
+MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in 8 billion parameters.
+It excels in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [19].
+InternVL2-8B features an InternViT-300M-448px vision encoder with 300 million parameters, paired with an internlm2_5-7b-chat language model comprising 7.7 billion parameters, totaling 8 billion parameters.
+Similarly, InternVL2-26B combines an InternViT-6B-448px-V1-5 vision encoder with 6 billion parameters and an internlm2-chat-20b language model containing 20 billion parameters, resulting in 26 billion parameters.
+B.3 SLO Criteria
+Tab.
+6 outlines the SLO criteria for TTFT and TPOT across various models and different numbers of images per request (#I/R).
+These criteria are empirically derived based on the characteristics of the underlying models, such as the computational complexity of the MME and LLM.
+We also consider what is realistically achievable by both our method and the baselines on a fixed number (8 GPUs) used in experiments.
+Across models, as #I/R increases, there is an approximately linear increase in the TTFT criteria due to the higher encoding load and additional tokens generated during the prefill stage.
+In contrast, TPOT requires only minor adjustments since it is not directly impacted by changes in #I/R.
+MiniCPMv 2.6 8B InternVL 8B InternVL 26B
+#I/R TTFT TPOT TTFT TPOT TTFT TPOT
+2 1.40 0.04 1.20 0.05 3.50 0.07
+4 2.60 0.04 2.40 0.06 7.05 0.08
+6 3.90 0.06 3.55 0.09 11.00 0.95
+8 5.10 0.06 5.00 0.18 15.00 0.15
+Table 6: Requirements for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
+B.4 Details for the Ablation Experiment of Offline Optimizer
+In this experiment, 100 User requests arrive in real-time with each request containing 6 images, and they all request for the MiniCPMv model.
+For simplicity, we limit the configuration space to a restricted search space, where each data-parallel (DP) worker uses the same batch size configuration, and both TP and PP are fixed to 1.
+We explore the batch size configurations for workers across different stages, the number of DP workers in each stage, and the decision to enable IRP.
+The config identified by the optimizer is as follows: batch sizes for the E, P, and D stages are 2, 1, and 128, respectively; the number of workers in each stage is 6, 1, and 1; and IRP is enabled.
+We uniformly random-sample 10 configs from the search space.
+To maintain the same computational cost, all 8 GPUs are utilized across different configs.
+Consequently, the search space is constrained to leverage 8 GPUs, which can be enforced through rejection sampling.
+The sample mean of the performance metric (goodput, TTFT, and TPOT) is presented in Table 4.
+Recall that the goodput metric is defined as the maximum request rate while maintaining an SLO attainment of no less than 90%.
+For a fair comparison, When evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is same as the goodput of the EPD system.
+B.5 NPU Implementation Details
+There are intrinsic differences between NPU and GPU architectures and processing characteristics.
+GPUs are general-purpose compute resources whereas NPUs have been specifically designed for neural networks.
+NPUs are optimized for FP16 calculations whereas GPUs are optimized for FP32 calculations.
+As stated above, CANN operators are used on Ascend NPU hardware in place of CUDA operators.
+These differences between CUDA and corresponding CANN operators, having been optimized for the hardware they are built upon, result in different latencies and trade-offs in inference as well as communication between NPU/GPU chips.
+This results in different encode/prefill/decode latencies which may be present different benefits from EPD and IRP.
+The EPD-NPU system was implemented in a similar manner to our EPD-GPU system as described in Appendix B.
+All EPD-NPU experiments utilize IRP.
+The EPD-NPU distributed serving system was developed and deployed on a cluster of 910B3 NPUs, each with 64GB of high bandwidth memory.
+Some primary differences between the two implementations include:
+  • Ascend-vLLM: EPD-NPU leverages Ascend-vLLM rather than vLLM which is an adaptation of vLLM on Ascend NPU hardware that utilizes the Compute Architecture for Neural Networks (CANN) software stack rather than CUDA.
+  • Container-based deployment: Each encode, prefill, and decode instance is deployed separately as an API for the scheduler engine to call.
+These instances run in data parallel mode.
+The scheduler engine simply sends a POST request to the corresponding API with some minimal request-specific information (e.g., request ID, sampling parameters, block IDs to pull) and the remaining implementation mirrors that of EPD-GPU.
+This container-based deployment allows our framework to be easily scalable and portable to the cloud, providing intriguing future works regarding efficient utilization and scheduling of resources for users.
+Figure 9: Left: The impact of tuning the number of encoding workers in the EPD method.
+The annotation ‘xEyP’ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P’ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests.
+Middle: The impact of the number of images per request in the workload.
+Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
+These differences in characteristics provide an opportunity to investigate the generalizability of our EPD framework and further understand the optimal setting/environment for disaggregating encode and prefill.
+For EPD-NPU, we use eight 910B3 NPUs and evaluate the system by running the controlled experimental setup (as described in the experiments section of the main paper) in an online setting.
+Specifically, the number of input tokens was 22, length of output tokens was 10 with early stopping was disabled.
+The SLO requirements were also selected in a similar manner as the GPU experiments.
+The requirements for InternVL2-8B with 8 images per request were 8.5 and 0.12 for TTFT and TPOT, respectively.
+Each server had 192 CPUs and 1.5TB of RAM.
+The CANN version was 7.6.
+The maximum model length was set to 27000, KV Cache utilization was set to 83% And the maximum number of multimedia data of 8 was imposed.
+EPD-NPU also used a version of Ascend-vLLM (version 0.6.3.post1) built upon a newer version of vLLM than the one used in EPD-GPU.
+Unless otherwise stated, the hyper-parameters are identical to those described in B.1.
+Appendix C Throughput in Offline Settings
+In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the E and P stages.
+Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput.
+This is an additional performance metric complementing the main body of the paper, measured in terms of requests completed per second and the system’s ability to handle high request volumes in offline settings.
+In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
+This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory.
+In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory.
+In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory.
+In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
 (a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
-Figure 4: SLO attainment for end-to-end inference across various models and image counts per request.
-Column (a) presents results for the MiniCPMv 2.6 model with 2, 4, 6, and 8 images per request (from top to bottom).
-Column (b) and (c) show results for InternVL2-8B and InternVL2-26B, respectively.
-The image resolution is fixed at 4K.
-As observed, EPD significantly outperforms all baselines across all configurations.
-EPD-IRP means that IRP is disabled in EPD system.
-5.1 SLO Attainment for End-to-End Generation
-In this experiment, we evaluate the end-to-end SLO performance of the proposed method (EPD), an ablation of our method without IRP (EPD-IRP) and existing state-of-the-art baselines, DistServe and vLLM.
-The setup involves 8 NVIDIA A100 GPUs serving 100 multimodal requests, arriving via a Poisson process with a fixed rate ($\lambda$) representing the number of requests per second.
-We tested three different models—MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B—under varying image counts per request All images were fixed at a 4K resolution of 4032x3024 pixels.
-The SLO criteria for each configuration, shown in Tab.
-1, establish latency upper bounds for both TTFT and TPOT, ensuring timely token generation and consistent decoding performance.
-The results for MiniCPMv 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right) are presented in Fig.
-4.
-Each row corresponds to a different image count per request, specifically 2, 4, 6, and 8, listed from top to bottom.
-The X-axis represents the per-GPU request rate (requests per second, req/s), while the Y-axis shows the SLO attainment percentage, defined as the proportion of requests meeting both TTFT and TPOT thresholds.
-The black dotted line indicates the 90% SLO attainment threshold, and the corresponding request rate represents the goodput for 90% SLO attainment.
-EPD-IRP demonstrates significantly higher SLO attainment compared to other baselines.
-At lower request rates, EPD-IRP achieves over 90% SLO attainment in most scenarios.
-This performance advantage is attributed to EPD-IRP’s ability to parallelize the computationally intensive image encoding step across multiple GPUs, mitigating the delay due to computationally-heavy encoding step.
-The second-best performer, EPD, achieves an SLO attainment close to 50% in many cases.
-However, the prior work, DistServe and vLLM, perform the worst, often attaining SLO percentages below 10%.
-These baselines suffer from interference between encoding and prefill stages, resulting in consistently poor SLO attainment.
-On the other hand, EPD can mitigate this interference due to disaggregation of encoding and prefill, however it lacks the parallelization ability of EPD-IRP.
-As the image count per request increases (from top to bottom rows), the SLO attainment for all baselines declines, particularly at higher request rates.
-This decline reflects the growing computational demands of processing multiple high-resolution images within a single request.
-In the most demanding scenario, with 16 images per request (bottom-most row), SLO attainment drops significantly for all baselines.
-However, EPD-IRP continues to maintain reasonable SLO attainment, especially at lower request rates, highlighting its robustness under heavy workloads.
-Finally, as we move from left to right across the plots, the change in model leads to progressively worse performance for all baselines.
-The InternVL models are prefill-heavy, with prefill stages taking substantially longer due to the generation of a higher number of image tokens.
-As a result, TTFT suffers from additional queuing delays at the prefill stage, especially at higher request rates.
-This effect becomes more pronounced when transitioning from the 8B (middle) to the 26B (right) model.
-In contrast, the MiniCPMv (left) model is optimized to generate fewer image tokens, making it encoding-heavy rather than prefill-heavy.
-Consequently, it avoids the queuing delays observed in the InternVL models and exhibits better latency characteristics overall.
-(a) MiniCPMv-2.6                                     (b) InternVL2-8B                                     (C) InternVL2-26B
-Figure 5: Distribution of TTFT (y-axis) for various #Images/ Request (x-axis) for the MiniCPMv 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right) models.
-Requests arrive following a Poisson distribution with a fixed request rate ($\lambda$).
-Note that, when comparing TTFT, vLLM is equivalent to DistServe, hence it is skipped.
-EPD-IRP means that IRP is disabled in EPD system.
-5.2 First Token Generation Latency
-In this experiment, we analyze the TTFT performance of the proposed method (EPD), an ablation of our method without IRP (EPD-IRP) and an existing state-of-the-art baseline, DistServe.
-Unlike prior analyses that focused on end-to-end inference, this experiment isolates the prefill phase of the computation, as the generation of the first token represents a critical latency bottleneck due to large number of prefill tokens in multimodel workloads.
-Note that in this setup the DistServe baseline is equivalent to vLLM since the decoding phase is not involved.
-Fig.
-5 shows the results for the three models: MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B, shown in the left, middle, and right plots, respectively.
-Each subplot presents box plots of TTFT distributions, with the x-axis representing the number of images per request and the y-axis showing the TTFT.
-The results are generated using 100 requests issued at a request rate of 0.25 for the MiniCPMv 2.6 and 0.08 for the InternVL2-8B and InternVL2-26B models.
-We limit the number of images per request (x-axis) on the basis of the respective model’s capacity.
-As shown in the figure, EPD-IRP significantly outperforms the two baselines, EPD and DistServe, across all three models.
-Specifically, the TTFT can be upto 71.9%, 32.8%, and 44.9% lower than the DistServe baseline for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
-This performance advantage stems from its intra-request parallelization, which enables the multimedia data within each request to be encoded in parallel, whereas the baselines process this data sequentially.
-Notably, EPD also demonstrates a slight performance improvement over DistServe, highlighting the independent benefits of disaggregating the encoding and prefill stages, even without full parallelization.
-In this case, the TTFT can be upto 9.8%, 4.9%, and 1.0% lower for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
-Further, we also observe that the margin of improvement becomes increasingly pronounced as the number of images per request grows.
-These results highlight the ability of EPD-IRP to efficiently handle the computational demands of multimodal workloads, especially as the number of input images and overall complexity increase.
-5.3 Memory Savings through Stage Disaggregation
-This section demonstrates the significant memory savings achieved by disaggregating the encoding and prefill stages in the EPD system.
-In existing systems, such as DistServe and vLLM, the encoding and prefill stages are co-located on the same GPUs, requiring each GPU to load both the MME and the LLM.
-This co-location results in higher memory consumption, as both models must be loaded simultaneously on the same GPUs.
-In contrast, the proposed disaggregated system optimizes memory usage by assigning the encoding stage to GPUs that load only the MME, while the prefill stage is handled by separate GPUs that load only the LLM.
-This separation significantly reduces memory requirements for both stages.
-The memory savings achieved through this disaggregation are substantial.
-Specifically, the encoding-stage GPUs experience memory reductions of approximately 95%, 96.2%, and 78.3% for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
-On the other hand, the prefill-stage GPUs achieve memory savings of about 5%, 3.7%, and 21.6% for the same models.
-In practice, the memory savings could be as high as 15x lower as encoding stage do not need to load the KV cache which is only required for prefill.
-These reductions in memory usage enable the EPD system to allocate resources more efficiently, allowing for the support of additional tasks and models, as demonstrated in the subsequent experiments.
-EPD Supports a Higher Number of Images per Request: Tab.
-2 compares the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B.
-The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and 80% of the available memory allocated for KV cache.
-For the MiniCPMv model, the EPD approach demonstrates a significant advantage in handling a higher number of images per request.
-At a resolution of 313x234, EPD supports up to 490 images per request, compared to just 77 for the aggregated DistServe system, resulting in a more than 6x improvement.
-At the 787x444 resolution, EPD can handle 165 images per request, which is a 6.3x increase over DistServe’s limit of 26.
-Even at the highest resolution (4032x3024), EPD supports 49 images per request, while DistServe is limited to only 7, demonstrating a 7x improvement.
-For the InternVL2-26B model, the improvements are equally striking.
-EPD supports up to 10x more images per request at both the 313x234 and 4032x3024 resolutions, and approximately 4x more at the 787x444 resolution, compared to DistServe.
-Interestingly, for the InternVL2-8B model, both EPD and DistServe support the same number of images per request across all resolutions (19).
-This is because the InternVL2-8B model model uses a relatively simple MME that is not memory-intensive, and the limit of 19 images is constrained by the maximum context length of the LLM rather than memory usage of the MME.
-Model Image Reso. DistServe EPD
-MiniCPMv 2.6 313,234 77 490
-787,444 26 165
-4032,3024 7 49
-InternVL2-8B 313,234 19 19
-787,444 19 19
-4032,3024 19 19
-InternVL2-26B 313,234 1 10
-787,444 11 45
-4032,3024 1 10
-Table 2: Comparison of maximum supported number of images/request.
-Batch size is fixed to 1.
-KV cache is allocated as 80% of the available free memory.
-Experiment done on A800 GPU with 82GB memory.
-Higher (italicized) is better.
-EPD Supports Higher Batch Sizes for Encoding and Prefill:
-Tab.
-3 compares the maximum supported batch sizes for the encoding (E) and prefill (P) stages across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B, at three different image resolutions.
-The number of images per request is fixed at 10, and each model uses a different patching strategy, leading to varying numbers of image patches.
-The last three columns show the maximum batch size that can be supported for each configuration without encountering out-of-memory (OOM) errors.
-For the EPD approach, batch sizes for encoding and prefill are listed separately, as these processes run independently on separate GPUs.
-In contrast, for the DistServe baseline, encoding and prefill share the same batch size since both stages are executed sequentially on a single GPU.
-The EPD approach significantly outperforms the DistServe baseline in terms of supported batch sizes.
-For instance, with the InternVL2-26B model at a resolution of 787x444, EPD supports a batch size of 22 for encoding, while the DistServe baseline can only handle a batch size of 1, representing a 22x improvement.
-On the other hand, with the MiniCPMv2.6 model at a resolution of 787x444, EPD supports a batch size of 29 vs 2 for the DistServe baseline, representing a 14.5x improvement for prefill.
-DistServe EPD
-Model Image Reso. #Patch E, P E P
-MiniCPMv 2.6 313,234 1 7 49 86
-787,444 3 2 16 29
-4032,3024 10 OOM 4 9
-InternVL2-8B 313,234 13 2 15 2
-787,444 3 9 67 10
-4032,3024 13 2 15 2
-InternVL2-26B 313,234 13 OOM 6 1
-787,444 3 1 22 4
-4032, 3024 13 OOM 6 1
-Table 3: Comparison of maximum supported E and P batch size.
-As seen, E, P batch sizes is much higher in EPD system in comparison to DistServe.
-Number of images/ request is fixed to 10.
-KV cache is allocated as 80% of the available free memory.
-Experiment done on one A800 GPU with 82GB memory.
-Higher (italicized) is better.
-EPD Supports Larger KV Cache Sizes: Tab.
-4 compares the maximum KV cache size that can be allocated in the disaggregated (EPD) system with that of the aggregated (DistServe or vLLM) system.
-In this experiment, the batch size is fixed at 1, and the image resolution is set to 4032×3024.
-We evaluate two LMM models, InternVL and MiniCPMv, with varying numbers of images per request.
-The results clearly demonstrate that EPD supports significantly larger KV cache sizes than DistServe.
-For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for Distserve, representing a 2.2x improvement.
-Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an Out of Memory (OOM) error, indicating that it cannot process 20 images even with a KV cache size of 0.
-Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
+Figure 10: Columns (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively.
+The top and bottom rows show results for 6 and 8 images per request, respectively.
+As expected, EPD maintains a significant performance advantage over all baselines, demonstrating robustness of the system as the image count increases.
+We use 8 A800 GPUs for this experiment.
+For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively.
+In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively.
+The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens.
+We tune the system’s hyperparameters as outlined in Sec.
+3.2.3 and evaluate their impact on throughput.
+The results are presented in Fig.
+9.
+The left plot illustrates the importance of selecting an appropriate GPU configuration.
+As described in Sec.
+3.2.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput.
+The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck.
+The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method.
+This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec.
+3.2.3 to automatically select an optimal batch size.
+Appendix D Additional Experiments
+D.1 SLO Attainment for End-to-End Generation for 6 and 8 images
+In this section, we show additional results from the experiment in Sec.
+4.1 pertaining to 6 and 8 images per request.
+As seen, the SLO attainment for EPD begins to decline as number of images rise, particularly at higher request rates.
+This decline reflects the increasing computational demand associated with processing multiple high-resolution images within a single request.
+Nevertheless, EPD continues to outperform all baselines, which struggle even at low request rates and fail to scale effectively under increased workloads.
 Model # Images/Req. DistServe EPD
-MiniCPMv 2.6 5 86% 99%
+MiniCPM-V 2.6 5 86% 99%
 10 74% 97%
@@
 Image resolution fixed to 4K.
-OOM refers to Out of Memory and OOCL refers to Out of Context Limit.
-Experiments are done on A800 GPU with 82GB memory.
 Higher (italicized) is better.
-Figure 6: Left: The impact of tuning the number of encoding workers in the EPD method.
-The annotation ‘xEyP‘ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P‘ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests.
-Middle: The impact of the number of images per request in the workload.
-Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
-5.4 Throughput in Offline Settings
-In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the encoding (E) and prefill (P) stages.
-Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput.
-In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
-This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory.
-In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory.
-In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory.
-In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
-We use 8 A800 GPUs for this experiment.
-For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively.
-In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively.
-The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens.
-We tune the system’s hyperparameters as outlined in Sec. 3.3.3 and evaluate their impact on throughput. The results are presented in Fig. 6.
-The left plot illustrates the importance of selecting an appropriate GPU configuration.
-As described in Sec.
-3.3.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput.
-The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck.
-The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method.
-This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec.
-3.3.3 to automatically select an optimal batch size.
-6 Conclusion
-In this paper, we present a novel approach to optimizing Large Multimodal Model (LMM) systems through the disaggregation of key processing stages.
-By separating the encoding, prefill, and decoding tasks into distinct stages, our system offers enhanced flexibility in resource allocation, enabling more efficient management of computational and memory resources.
-This disaggregation, combined with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage.
-Our results demonstrate significant improvements in both throughput and memory efficiency, particularly in heterogeneous environments where resources are limited.
-By efficiently parallelizing the encoding and prefill stages, we are able to mitigate common bottlenecks and improve the scalability of multimodal inference tasks.
-Furthermore, the optimized memory management strategies ensure that the system can scale to handle larger models and more complex multimodal inputs, which are increasingly common in real-world applications.
-The approach presented in this work represents a significant step forward in the development of scalable and efficient systems for multimodal AI, offering a foundation for future research and practical applications in areas such as natural language processing, computer vision, and cross-modal understanding.
-In summary, our system provides a robust solution for improving the performance of LMMs, making them more efficient and adaptable for deployment in real-world, resource-constrained environments.
+(a) GPU                        (b) NPU
+Figure 11: Breaking down the latency of encode and prefill for InternVL2-8B across different number of images per requests.
+We compare the difference between GPU (left) and NPU (right) where light green is encode and light blue is prefill.
+D.2 EPD Supports Larger KV Cache Sizes:
+We compare the maximum KV cache size that can be allocated across systems.
+The batch size is fixed at 1, and the image resolution is set to 4032$\times$3024.
+As shown in Tab.
+7, EPD supports significantly larger KV cache sizes than DistServe.
+For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for DistServe, representing a 2.2$\times$ improvement.
+Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an OOM error, indicating that it cannot process 20 images even with a KV cache size of 0.
+Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
+D.3 Heavier Encode in NPUs
+We compare the encode-to-prefill stage latencies between NPUs and GPUs.
+Figure  11 show the results for the InternVL2-8B model, across various I/R values.
+As seen, we notice higher ($\approx 10-20\%$) encode-to-prefill latency ratios for NPUs in comparison to GPUs, indicating the need for a different optimized configuration for NPU.
+This trend was consistent across different large multimodal models (LMMs), indicating that NPUs tend to spend a greater proportion of time in the encoding phase compared to GPUs.

### v3

@@
 Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead.
-This step negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput (E2ETP).
+This step negatively affects key Service Level Objectives (SLOs), such as time to first token (TTFT) and time per output token (TPOT).
 We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources.
-Unlike current systems, which bundle encoding and prefill together, our approach decouple these steps unlocking new opportunities and optimizations.
-These include a new mechanism to cache multimedia tokens for efficient transfer, a novel way to parallelize encoding load within a request, a module to find the optimal resource allocation for disaggregated serving, and a novel role switching method to handle changing workload characteristics.
-Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ less utilization), batch sizes (up to 22$\times$ larger), 10$\times$ more images/request, and 2.2$\times$ larger KV caches.
-Further, it leads to significant improvements in latency metrics (TTFT up to 71% reduction) and end-to-end throughput (up to 57% reduction), compared to systems that do not disaggregate.
+Unlike current systems, which bundle encoding and prefill together, our approach decouples these steps, unlocking new opportunities and optimizations.
+These include a mechanism to cache multimedia tokens for efficient transfer, a novel way to parallelize the encoding load within a request, a module for optimal resource allocation for disaggregated serving, and a novel role-switching method to handle changing workload characteristics.
+Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ lower peak memory utilization), batch sizes (up to 22$\times$ larger), 10$\times$ more images per request, and 2.2$\times$ larger KV caches.
+Furthermore, it leads to significant improvements in SLO attainment (up to 90–100% improvement) and TTFT (up to 71% reduction), compared to systems that do not disaggregate.
+The code is available at https://github.com/vbdi/epdserve.
+Efficient inference, Disaggregated inference, Multimodal modal inference, LLM inference
 1 Introduction
-Large Language Models (LLMs) have revolutionized language understanding and reasoning, achieving superhuman performance across a variety of tasks [1, 4].
-Recently, the scope of these models has expanded to include multiple modalities, such as images, audio, and videos, resulting in the advent of Large Multimodal Models (LMMs) [19, 11, 5].
-LMMs enable users to interact with diverse data types, such as posing questions about visual scenes or analyzing audio clips, thereby unlocking novel applications across fields like healthcare, autonomous systems, and creative industries.
-Figure 1: Above: Aggregated system, where E and LLM are coupled and must run on the same GPU.
-Below: Disaggregated system, where E runs on the 1st GPUs and LLM runs on the rest of three GPUs.
-Multimodal encoder and LLM are denoted by E and LLM, respectively.
+Large Language Models (LLMs) have revolutionized language understanding and reasoning, achieving superhuman performance on a variety of tasks (Achiam et al., 2023; Chang et al., 2024).
+Recently, the scope of these models has expanded to include multiple modalities, such as images, audio, and videos, leading to the emergence of Large Multimodal Models (LMMs) (Yao et al., 2024; Liu et al., 2023; Chen et al., 2024).
+LMMs enable users to interact with diverse data types, such as posing questions about visual scenes or analyzing audio clips, thereby unlocking novel applications across fields like healthcare, autonomous systems, and creative industries (Luo et al., 2025).
+Figure 1: Aggregated (top) vs.
+disaggregated (bottom) system architectures.
+In the aggregated setup, the encoder (E) and LLM share the same GPUs, leading to interference between encode and prefill stages (e.g., LLM-4 delays E5).
+Disaggregation isolates these stages across GPUs, reducing contention and improving utilization.
+This separation enables better performance under multimodal workloads, addressing key limitations of existing systems.
 However, serving LMMs in an efficient manner presents unique challenges.
-Meeting strict Service Level Objectives (SLOs) such as end-to-end throughput (E2ETP), time to first token (TTFT), and time per output token (TPOT) becomes increasingly difficult given the additional computational and memory demands introduced by processing multimodal data [12].
+Meeting strict Service Level Objectives (SLOs), such as time to first token (TTFT) and time per output token (TPOT), becomes increasingly difficult given the added computational and memory demands of processing multimodal data (Alvar et al., 2025; Liu et al., 2024).
 Unlike LLMs, where inference involves prefill and decoding stages, LMMs require an additional encoding stage to process raw multimodal inputs (e.g., images or videos) into tokenized representations.
-This stage is computationally intensive, especially for high-resolution or complex multimodal inputs, and often produces a substantial number of additional tokens [17].
-The resulting increase in tokens inflates resource consumption and leads to quadratic growth in prefill-stage compute demands, adversely impacting SLO attainment.
-Disaggregating the prefill stage from the decode stage has emerged as a well-studied solution in the literature to improve inference efficiency for LLMs [21, 15, 14, 7, 6].
+This stage is computationally intensive, especially for high-resolution or complex multimodal inputs, and often produces a substantial number of additional tokens (Wu et al., 2023).
+The resulting token inflation increases resource consumption and leads to quadratic growth in prefill-stage compute demands, adversely impacting SLO attainment.
+Disaggregating the prefill stage from the decode stage has emerged as a well-studied solution for improving LLM inference efficiency (Zhong et al., 2024; Qin et al., 2024; Patel et al., 2024; Jin et al., 2024; Hu et al., 2024).
 By assigning separate resources to each stage, prefill-decode disaggregation enables independent optimization of batching, scheduling, and resource allocation strategies, significantly enhancing system throughput and memory utilization.
-However, these techniques fail to address the challenges of LMM deployment, where the addition of an encoding stage fundamentally alters the resource dynamics.
-The encoding stage, with its high computational and memory overhead, introduces token inflation and dependency bottlenecks that ripple across subsequent stages, necessitating a rethinking of disaggregation strategies to accommodate multimodal workloads.
-Figure 2: Impact of disaggregation on batch size and number of images/ request for the MiniCPM-V 2.6 model
-These challenges of LLM deployment present opportunities for optimization that current serving systems do not consider.
-Currently, the encode and prefill stages are aggregated into a single monolithic and synchronous step executed on the same set of GPUs.
-Sequential execution of encoding and prefilling introduces interference.
-This interference is demonstrated graphically in Figure 1.
-In the aggregated approach (Figure 1 top), the prefill step ($LLM^{4}$) interferes with the encode-heavy request ($E^{5}$).
-As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, revealing that current solutions are inadequate for LMM workloads.
-By disaggregating encode from prefill, the system may be optimized to reduce interferences for specific workloads.
-From another perspective, simultaneous loading of both the Multimodal encoder (MME) and LLM onto the same GPUs in aggregated setups restricts memory for processing multiple high resolution images or larger batch sizes.
-In preliminary investigations, by removing the LLM and associated KV cache, we observe a significant increase in maximum batch sizes and supported number of images/ request (Figure 2).
-Thus, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encode, prefill, and decode stages, assigning each with dedicated resources to operate independently.
-This disaggregation enables customized batching, parallelization, and scheduling strategies for each stage, optimizing resource utilization while reducing contention.
-As a result, EPD disaggregation achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT, TPOT, and E2ETP.
+However, these techniques fall short in addressing LMM-specific challenges, as the addition of an encoding stage fundamentally changes the resource dynamics.
+The encoding stage adds significant compute and memory overhead, inflates tokens, and creates dependencies that affect later stages, requiring a fresh look at disaggregation strategies for multimodal workloads.
+Figure 2: Impact of disaggregation on supported batch size and number of images per request for the MiniCPM-V 2.6 model.
+Removing the LLM from the GPU significantly increases capacity, enabling larger batches and higher-resolution inputs.
+This demonstrates the memory efficiency benefits of disaggregation.
+These challenges present optimization opportunities that current serving systems do not exploit.
+Presently, the encode and prefill stages are aggregated into a single monolithic and synchronous step executed on the same set of GPUs.
+Sequential execution of encoding and prefilling introduces interference, as demonstrated in Figure 1.
+In the aggregated setup (top), the prefill step ($LLM^{4}$) interferes with the encode-heavy request ($E^{5}$).
+As a result, such aggregation leads to suboptimal resource utilization and degraded SLO performance, revealing the inadequacy of current solutions for LMM workloads.
+Disaggregating encoding from prefill enables the system to reduce such interference under specific workloads.
+From another perspective, simultaneously loading both the multimodal encoder (MME) and the LLM onto the same GPUs in aggregated setups restricts the available memory for processing multiple high-resolution images or supporting larger batch sizes.
+In preliminary investigations, removing the LLM and its associated KV cache significantly increases the maximum batch size and the number of images per request (Figure 2).
+Thus, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encode, prefill, and decode stages, assigning each stage dedicated resources to operate independently.
+This design enables customized strategies for batching, parallelization, and scheduling at each stage, optimizing resource utilization while reducing contention.
+As a result, EPD achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT and TPOT.
 The major contributions of this work are as follows:
-  • We propose an efficient system for LMM inference that relies on the novel idea of disaggregating encoding and prefill stages.
-Our approach addresses key challenges of inter-stage communication, efficient parallelization, resource allocation, and performance optimization within this framework.
-  • We propose intra-request parallelization (IPR) that shards a request into independent encoding jobs which can be executed in parallel, significantly reducing first token latency.
+  • We propose an efficient system for LMM inference that introduces the novel idea of disaggregating the encoding and prefill stages.
+Our approach addresses key challenges in inter-stage communication, efficient parallelization, resource allocation, and performance optimization within this framework.
+  • We introduce intra-request parallelization (IRP), which shards a request into independent encoding jobs that can be executed in parallel, significantly reducing first-token latency.
   • We formulate the resource allocation problem as an optimization over batch sizes, scheduling strategies, and parallelization approaches for each pipeline stage.
-Further, we provide a simple black box optimizer that uses workload samples to approximate the optimal config for a given workload.
-  • We develop dynamic role switching capability that monitors the system for any bottlenecks and allows efficient reallocation of resources between E, P and D stages by efficiently switching ($\sim$0.7s) the role of an instance among these stages.
-This ensures system can respond to any changes in workload requirements during online serving.
-  • We conduct evaluations on various popular LMM models, namely MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B using real-world and synthetic workloads.
-The results demonstrate the superiority of our method across various aspects.
-For instance, our approach requires up to 15x less memory, facilitating up to 22x larger batch sizes, 10x more images/request, and 2.2x larger KV caches.
-Further, our method significantly improves performance in terms of SLO attainment (up to 90–100% more) and TTFT latency (up to 71% reduction).
+Furthermore, we provide a simple black-box optimizer that uses workload samples to approximate the optimal configuration for a given workload.
+  • We develop a dynamic role-switching capability that monitors the system for bottlenecks and enables flexible reallocation of resources across the E, P, and D stages by switching an instance’s role between stages.
+This ensures the system can respond effectively to changes in workload requirements during online serving.
+  • We conduct evaluations on several popular LMM models—MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B—using both real-world and synthetic workloads.
+Results demonstrate the superiority of our approach: up to 15$\times$ lower memory usage, 22$\times$ larger batch sizes, 10$\times$ more images per request, up to a 90–100% improvement in SLO attainment, and up to 71% lower TTFT.
 2 Related Work
-In this section, we describe feasible approaches in the literature for multi-modality model serving and their limitations and then review the use of the disaggregation technique.
-Multi-modality Model Serving.
-There are two ways to enable a multi-modality model serving system: (1) adopting and extending existing LLM serving systems, such as vLLM [9], SARATHI [2], and Orca [20]; and (2) leveraging the open-sourced code from recent works of improving multi-modality model inference.
-The first way fails to address the encoding bottleneck inherent in LMMs, limiting their applicability to multimodal workloads, particularly when handling high-volume and high-resolution multi-media data.
-In the second way, recent advancements such as KV cache eviction [10, 13] and compression [8] focus on specific features and limited scenarios.
-For instance, Inf-MLLM [13] enables efficient streaming inference for LMMs on a single GPU, targeting resource-constrained scenarios.
-While their implementations are feasible for serving LMMs, they fail to meet user SLOs in cloud serving scenarios.
+In this section, we describe existing approaches in the literature for multimodal model serving and their limitations, and then we review the use of disaggregation techniques.
+Multimodal Model Serving.
+There are two ways to enable a multimodal model serving system: (1) adopting and extending existing LLM serving systems, such as vLLM (Kwon et al., 2023), SARATHI (Agrawal et al., 2023), and Orca (Yu et al., 2022); and (2) leveraging open-source code from recent works on improving multimodal model inference.
+The first approach fails to address the encoding bottleneck inherent in LMMs, limiting their applicability to multimodal workloads, particularly when handling high-volume and high-resolution multimedia data.
+In the second approach, recent advancements such as KV cache eviction (Li et al., 2024; Ning et al., 2024) and compression (Kang et al., 2024) focus on specific features and limited scenarios.
+For instance, Inf-MLLM (Ning et al., 2024) enables efficient streaming inference for LMMs on a single GPU, targeting resource-constrained scenarios.
+While these implementations are feasible for serving LMMs, they fail to meet user SLOs in cloud serving scenarios.
 To the best of our knowledge, we are the first to propose an LMM serving system that leverages disaggregation and a series of integrated techniques to enable better resource allocation and improved SLOs.
@@
 Disaggregated serving has emerged as a promising technique in large model serving.
-By decoupling the prefill and decode stages, systems like SplitWise [14], DistServe [21], and DéjàVu [16] mitigate interference between these phases, enabling finer control over Time to First Token (TTFT) and Time Per Output Token (TPOT).
+By decoupling the prefill and decode stages, systems like SplitWise (Patel et al., 2024), DistServe (Zhong et al., 2024), and DéjàVu (Strati et al., 2024) mitigate interference between these phases, enabling finer control over TTFT and TPOT.
 However, these systems primarily target LLMs and overlook the encoding step required for LMMs, which is tightly coupled with prefill.
-Recent work, such as Mooncake [15] and PD-Serve [7], has extended disaggregation to include KV cache management and other system-level optimizations.
+Recent work, such as Mooncake (Qin et al., 2024) and PD-Serve (Jin et al., 2024), have extended disaggregation to include KV cache management and other system-level optimizations.
 Despite these advancements, they remain limited in their applicability to LMMs.
 Disaggregating the encoding phase from prefill could unlock new opportunities for optimizing LMM serving.
-For instance, when handling requests with long videos, encoding the frames could be executed in parallel, reducing head-of-line blocking and improving overall system efficiency.
-This approach would enable more flexible batching and scheduling strategies, ultimately achieving better latency and throughput for multimodal workloads.
+For instance, when handling requests with long videos, the frames could be encoded in parallel, reducing head-of-line blocking and improving overall system efficiency.
+This approach would enable more flexible batching and scheduling strategies, achieving better latency and throughput for multimodal workloads.
 3 Method
-In this section, first, we describe the EPD disaggregation framework (Sec.
-3.1), followed by a detailed explanation of the system design and optimizations (Sec.
-3.2), that includes the asynchronous token transfer, intra-request parallel, optimized resource allocation, and dynamic role switching.
+In this section, we first describe the EPD disaggregation framework (Section 3.1), followed by a detailed explanation of the system design and optimizations (Section 3.2), including asynchronous token transfer, intra-request parallelism, optimized resource allocation, and dynamic role switching.
 Figure 3: The inference pipeline of EPD Disaggregation.
 3.1 EPD Disaggregation
-As shown in Figure 3, disaggregating an LMM system involves breaking the inference process into multiple stages: encoding, prefill, and decode.
+As shown in Figure 3, disaggregating an LMM system involves dividing the inference process into three stages: encoding, prefill, and decode.
 Transitions between these stages—EP-migration and PD-migration—handle the transfer of data from encoding to prefill and from prefill to decode, respectively.
-Formally, we denote the input text prompt as $i_{p}$, multimodal data as $i_{m}$, and the text output as $o$, and provide details for each step and transition as follows:
-Encoding (E): The multimodal input $i_{m}$ is processed by the MME $E$, converting it into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs.
-These tokens form a high-dimensional multimodal embedding for the next pipeline stage.
-This process can be denoted as $v_{t}^{e}=E(i_{m})$.
-EP-migration: Once the $E$ stage completes processing, the generated multimodal tokens $v_{t}^{e}$ are transferred to the prefill stage $P$.
-This transfer is managed by the EP-migration function $\psi_{EP}$, which handles the migration of tokens from the encoding worker’s cache to the prefill worker’s cache.
-This process is expressed as $v_{t}^{p}=\psi_{EP}(v_{t}^{e})$.
-Here, $v_{t}^{p}$ represents the tokens in the prefill stage after migration.
-Prefill (P): The multimodal tokens $v_{t}^{p}$ arrive at the Prefill stage $P$, along with the text prompt $i_{p}$.
-In this stage, the system initializes the Key-Value (KV) cache $kv_{1}^{p}$ and generates the first output token $o_{1}^{p}$.
-The KV cache stores context and intermediate states, which are necessary for the subsequent decoding stage.
-The process is denoted by $kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$.
-PD-migration: After the prefill stage generates the initial token $o_{1}^{p}$ and updates the KV cache to $kv_{1}^{p}$, the data must be transferred to the decode stage $D$.
-This transfer is managed by the PD-migration function $\psi_{PD}$, which ensures that the KV cache and the generated token are correctly handed off to the decode worker.
-This is denoted as $kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$.
-Decode (D): The decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the token generated in the previous iteration $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$.
-This autoregressive process continues until the desired output sequence is generated, as denoted by $kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$.
+We denote the input text prompt as $i_{p}$, multimodal data as $i_{m}$, and the output text as $o$.
+The steps are as follows:
+Encoding (E): The multimodal input $i_{m}$ is processed by the multimodal encoder $E$, converting it into multimodal tokens $v_{t}^{e}=E(i_{m})$, which form a high-dimensional embedding for the next pipeline stage.
+EP-migration: Once encoding completes, the generated tokens are transferred to the prefill stage $P$ via the EP-migration function $\psi_{EP}$, such that $v_{t}^{p}=\psi_{EP}(v_{t}^{e})$.
+Prefill (P): The prefill stage processes $v_{t}^{p}$ and the text prompt $i_{p}$ to produce the initial KV cache and first output token: $kv_{1}^{p},o_{1}^{p}=P(v_{t}^{p},i_{p})$.
+PD-migration: The KV cache and token are passed to the decode stage via the PD-migration function $\psi_{PD}$, where $kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$.
+Decode (D): Decoding proceeds autoregressively, generating the next token $o_{t+1}^{d}$ and updating the cache: $kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$.
+This autoregressive process continues until the output sequence is fully generated.
 Figure 4: System architecture of the proposed EPD Disaggregated Inference.
@@
 Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on ensuring smooth token transfers between stages (Section 3.2.1), reducing latency (Section 3.2.2), managing resources effectively (Section 3.2.3), and adapting to changing workloads (Section 3.2.4).
-The ablation of these features are shown in Sec.
-4.4.
+The ablation of these features is shown in Section 4.4.
 3.2.1 Asynchronous Token Transfer
@@
 3.2.2 Intra-Request Parallel (IRP)
-Multimodal requests often include multiple high-resolution images in many practical scenarions like autonomous driving and video question answering.
-In modern LMMs, these images are further converted into a large number of patches that needs to be processed through a computationally-heavy MME.
+Multimodal requests often include multiple high-resolution images in many practical scenarios like autonomous driving and video question answering.
+In modern LMMs, these images are further converted into a large number of patches that need to be processed through a computationally heavy MME.
 This significantly increases the computation load, making the encoding process a major bottleneck.
@@
 Let $f(\cdot)$ represent the system’s performance metric (e.g., goodput, see Section 4).
-Since $f(\cdot)$ is treated as a black-box function with an unknown internal mechanism, we rely on a simulator—extended from DistServe [21]—to evaluate performance metrics efficiently.
+Since $f(\cdot)$ is treated as a black-box function with an unknown internal mechanism, we rely on a simulator—extended from DistServe (Zhong et al., 2024)—to evaluate performance metrics efficiently.
 The objective is to maximize performance while minimizing GPU usage, which inherently reduces pipeline inefficiencies (e.g., idle time) and improves resource utilization.
@@
 $\max_{(\mathbf{p},\mathbf{b},\mathbf{s})\in\mathcal{X}}f(\mathbf{p},\mathbf{b},\mathbf{s})-\beta cost(\mathbf{p})$ (1)
-Here, $\mathcal{X}$ is the search space for system configs, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$ (See Appendix A for details).
-We use Bayesian optimization [3] to solve Problem 1.
+Here, $\mathcal{X}$ is the search space for system configs, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$ (See Appendix D for details).
+We use Bayesian optimization (Calvo et al., 2019) to solve Problem 1.
 3.2.4 Dynamic Role Switching
@@
 In this section, we analyze and compare the performance of the proposed EPD disaggregation method against various baselines.
-We start with an end-to-end generation performance analysis in Sec.
-4.1, followed by an examination of first token latency in Sec.
-4.2.
-We then evaluate the memory savings achieved by EPD in Sec.
-4.3.
-Finally, an ablation analysis of key system components is presented in Sec.
-4.4.
-Additional experiments are detailed in the Appendix, including 1) a throughput comparison in offline settings (Appendix C); 2) an extension of our framework to Neural Processing Units (NPUs) demonstrating a 10% performance improvement (LABEL:sec:appendix-npu), and 3) additional SLO and memory experiments (Appendix D).
+We start with an end-to-end generation performance analysis in Section 4.1, followed by an examination of the first token latency in Section 4.2.
+We then evaluate the memory savings achieved by EPD in Section 4.3.
+Next, we present an ablation analysis of the key system components in Section 4.4.
+Finally, we present an extension of our framework to Neural Processing Units (NPUs) in Section 4.5.
+Additional experiments are detailed in the Appendix, including 1) a throughput comparison for offline and heterogeneous settings in Section A.3, 2) additional analysis of the implementation on NPUs in Appendix F, 3) an experiment extending our framework to the audio modality in Section A.1, and 4) additional SLO and memory experiments in Appendix A.
 (a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
-Figure 5: SLO attainment for end-to-end inference across various models and image counts per request.
-Columns (a), (b), and (c) correspond to the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively.
-The top and bottom rows are with 2 and 4 images per request, respectively.
-As observed, EPD significantly outperforms all baselines across all configurations.
-Baselines: We compared the proposed EPD method against two popular baselines: DistServe [21] and vLLM [9].
-The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on separate GPUs.
-Since DistServe was originally designed for LLMs, we extended it to support LMMs by enabling multimodal data processing and modifying the block manager to accommodate multimodal tokens.
+Figure 5: SLO attainment ($\uparrow$) for end-to-end inference across multiple models and image counts per request.
+Subfigures (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively.
+The top and bottom rows show results for 2 and 4 images per request.
+EPD consistently outperforms all baselines across configurations.
+Baselines: We compared the proposed EPD method against two popular baselines: DistServe (Zhong et al., 2024) and vLLM (Kwon et al., 2023).
+The DistServe baseline implements the prefill-decode (PD) disaggregation approach, in which the prefill and encoding phases are executed on one set of GPUs, while the decode phase is disaggregated on separate GPUs.
+Since DistServe was originally designed for LLMs, we extended it to support LMMs by enabling multimodal data processing and modifying its block manager to accommodate multimodal tokens.
 The vLLM baseline adopts a monolithic architecture, where all three stages run on the same set of GPUs.
-Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [19], InternVL2-8B, and InternVL2-26B [5].
+Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 (Yao et al., 2024), InternVL2-8B, and InternVL2-26B (Chen et al., 2024).
 These LMMs are renowned for their advanced capabilities in processing and understanding multimodal data.
-Detailed descriptions of the LMMs and their sizes can be found in Appendix B.2.
-Datasets: To evaluate performance across diverse scenarios, we conducted experiments using two datasets.
-The first was a synthetic workload, that allowed setting custom parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings.
-Unless otherwise specified in our experiments, the input prompt length was set to 22 tokens.
-The second dataset was NextQA [18], a benchmark video question-answering dataset.
-Unlike synthetic workload, NextQA features human-annotated questions and answers, providing a more realistic reflection of real-world video request distributions.
-Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The related performance metrics are listed below.
+Detailed descriptions of the LMMs and their sizes can be found in Appendix E.2.
+Datasets: To evaluate performance across diverse scenarios, we use three datasets: synthetic workload, NextQA, and Video-MME.
+The synthetic workload enables configurable parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings.
+Unless otherwise noted, the input prompt length is set to 22 tokens.
+NextQA (Xiao et al., 2021), a benchmark video question-answering dataset, features human-annotated questions and answers, offering a more realistic reflection of real-world video request distributions compared to the synthetic workload.
+Video-MME (Fu et al., 2024) is a multimodal evaluation dataset designed for assessing large LMMs on video understanding tasks.
+It contains multiple-choice video question answering items that span diverse video lengths, topics, and reasoning types.
+In contrast to NextQA, which emphasizes open-ended video-text generation, Video-MME focuses strictly on multiple-choice QA and covers a broader spectrum of video durations and formats.
+Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The performance metrics are as follows:
   • TTFT: The time from the submission of a request to the system until the first token is received by the user.
-  • TPOT: The average time interval between each output token (excluding the first token).
+  • TPOT: The average time interval between consecutive output tokens (excluding the first token).
   • SLO Attainment: The percentage of requests that meet predefined SLOs, such as TTFT and TPOT requirements.
-  • Goodput: The hightest request rate at which 90% or more SLO attainment is achieved.
-For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory.
-Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes.
-(See section 4.3)
+  • Goodput: The highest request rate at which 90% or more SLO attainment is achieved.
+For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory in each approach.
+Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes (see Section 4.3).
 4.1 SLO Attainment for End-to-End Generation
-In this experiment, we evaluate the goodput of EPD and the baselines in an online setting, where 100 multimodal requests arrive following a Poisson process with rate $\lambda$.
+In this experiment, we evaluate the goodput of EPD and the baselines in an online setting where 100 multimodal requests arrive following a Poisson process with rate $\lambda$.
 The length of output tokens is fixed to 10.
 We test requests with 2 and 4 images, each at a resolution of $4032\times 3024$.
-The details of the corresponding SLO criteria are discussed in Sec.
-B.3.
-The results for three models (columns) are presented in Fig.
-5.
-The X-axis represents the per-GPU request rate ($\lambda$/#GPUs), and the Y-axis indicates the percentage of requests meeting both TTFT and TPOT requirements, with a 90% threshold denoted by a black dotted line.
+The details of the corresponding SLO criteria are discussed in Section E.3.
+The results for three models (columns) are presented in Figure 5.
+The X-axis represents the overall request rate ($\lambda$), and the Y-axis indicates the percentage of requests meeting both TTFT and TPOT requirements, with a 90% threshold denoted by a black dotted line.
 (a) MiniCPMv-2.6                  (b) InternVL2-8B                     (c) InternVL2-26B
-Figure 6: Distribution of TTFT (Y-axis) for various #Images/ Request (X-axis) for the MiniCPM-V 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right)
+Figure 6: Distribution of TTFT (Y-axis) across varying numbers of images per request (X-axis) for (a) MiniCPM-V 2.6, (b) InternVL2-8B, and (c) InternVL2-26B.
+Each plot illustrates the latency behavior under increasing input sizes.
 EPD outperforms all baselines, achieving over 90% SLO attainment at lower request rates.
 This is due to its ability to parallelize the computationally intensive image encoding step across multiple GPUs.
-DistServe and vLLM often maintain less than 10% attainment due to interference.
-Meanwhile, comparing the first and second rows for varying image counts per request (2 and 4), we observe that while more images per request increase workload, EPD maintains reasonable performance, whereas baseline SLO attainment drops significantly.
-Lastly, InternVL, being prefill-heavy, experiences queuing delays due to the higher number of image tokens, particularly at higher request rates.
-This effect exacerbates performance degradation from the 8B to the 26B model.
-In contrast, MiniCPM-V, optimized to generate fewer image tokens, avoids these delays and achieves better overall latency.
-Further, we provide more results with 6 and 8 images in Sec.
-D.1.
-Next, we repeat the experiment using the non-synthetic video question-answering dataset, NextQA [18].
-For this, we randomly sampled 100 examples, with input text token lengths ranging from 4 to 21 (average: 11.42) and output token lengths ranging from 1 to 7 (average: 2.75).
+DistServe and vLLM often maintain less than 10% SLO attainment due to interference.
+Meanwhile, comparing the first and second rows for varying image counts per request (2 and 4), we observe that, while more images per request increase the workload, EPD maintains reasonable performance, whereas baselines’ SLO attainment drops significantly.
+Lastly, InternVL, which is prefill-heavy, experiences queuing delays due to the higher number of image tokens, particularly at higher request rates.
+This effect further exacerbates the performance degradation from the 8B to the 26B model.
+In contrast, MiniCPM-V, optimized to generate fewer image tokens, avoids these delays and achieves lower overall latency.
+Further results with 6 and 8 images are provided in Section A.4.
+Next, we repeat the experiment using the non-synthetic video question-answering dataset, NextQA (Xiao et al., 2021).
+To do so, we randomly sampled 100 examples, with input text token lengths ranging from 4 to 21 (average: 11.42) and output token lengths ranging from 1 to 7 (average: 2.75).
 Each video request was represented by 8 uniformly sampled frames, and we used MiniCPM-V 2.6 in the experiment, adhering to the SLO criteria of TTFT = 5.60 and TPOT = 0.06.
-As illustrated in Fig.
-7, EPD is the only framework achieving 90% SLO attainment at low request rates, demonstrating its superior effectiveness in handling diverse, real-world-like workloads compared to DistServe and vLLM.
-Figure 7: SLO attainment for end-to-end inference on the NextQA dataset. As seen, EPD outperforms other baselines by a significant margin.
+As illustrated in Figure 7, EPD is the only framework achieving 90% SLO attainment at low request rates, demonstrating its superior ability to handle real-world workloads compared to DistServe and vLLM.
+Figure 7: SLO attainment ($\uparrow$) versus request rate on the NextQA dataset using the MiniCPM-V 2.6 model. EPD consistently achieves higher SLO attainment compared to all baselines.
+Finally, we conduct the same experiment using the Video-MME (Fu et al., 2024) dataset.
+We evaluate SLO attainment, defined as TTFT $\leq$ 3.1s and TPOT $\leq$ 0.025s, on 100 randomly sampled Video-MME examples using MiniCPM-V 2.6.
+Each video is represented by 64 uniformly sampled frames, following the MiniCPM frame configuration reported on the Video-MME leaderboard.
+The results are shown in Figure 8.
+As seen, EPD consistently outperforms vLLM and DistServe across all rates, demonstrating a strong generalization to temporal multimodal workloads, such as videos.
+Figure 8: SLO attainment ($\uparrow$) versus request rate on the Video-MME dataset using the MiniCPM-V 2.6 model. EPD significantly outperforms competing baselines across all request rates.
+Method #Frames 8 16 32 64
+vLLM 0.42 0.82 1.59 3.11
+DistServe 0.42 0.81 1.54 3.08
+EPD (ours) 0.24 0.30 0.49 1.00
+Table 1: Mean TTFT latency (in seconds) ($\downarrow$) for varying video lengths at a fixed request rate of 1 request/sec.
+Results are averaged over 100 Video-MME samples.
+EPD achieves the lowest latency across all video lengths.
 4.2 First Token Generation Latency
-Multimodal requests often place a heavy load on both the E and P phases, making the first token generation a key latency bottleneck.
+Multimodal requests often impose a heavy load on both the E and P phases, making the first token generation a key latency bottleneck.
 Therefore, in this experiment, we analyze the first token latency for various baselines.
-The results presented in Fig.
-6 shows box plots of TTFT distributions for three different models.
-Note that, as the decoding phase is excluded, the vLLM baseline is equivalent to DistServe and is thus skipped.
+The results presented in Figure 6 show box plots of TTFT distributions for three different models.
+Note that, as the decoding phase is excluded, the vLLM baseline is equivalent to DistServe and is thus omitted.
 Requests are generated according to a Poisson distribution with a fixed request rate $\lambda$.
-Specifically, $\lambda=0.25$ for the MiniCPM-V 2.6, and $\lambda=0.08$ for both the InternVL2-8B and InternVL2-26B.
-Thanks to intra-request parallelization, EPD significantly outperforms both baselines, EPD and DistServe.
+Specifically, $\lambda=0.25$ for MiniCPM-V 2.6 and $\lambda=0.08$ for both InternVL2-8B and InternVL2-26B.
+Thanks to intra-request parallelization, EPD significantly outperforms both vLLM and DistServe.
 Specifically, TTFT is reduced by up to 71.9%, 32.8%, and 44.9% compared to the DistServe baseline for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
-The performance gain grows with the number of images per request, demonstrating EPD’s efficiency in handling heavy multimodal workloads.
+Additionally, we conducted a TTFT comparison on the Video-MME (Fu et al., 2024) dataset for various baselines.
+The results are presented in Table 1.
+As seen, EPD achieves a latency reduction of up to 68.2% over DistServe and up to 69.2% over vLLM.
+Moreover, the performance gap widens with increasing video length—for instance, at 8 frames, EPD reduces latency by 42.9%, and at 64 frames, the reduction reaches 67.5% relative to DistServe.
+These results highlight EPD’s superior scalability and robustness under increasingly demanding video processing workloads.
 4.3 Memory Savings through Stage Disaggregation
-In this section we analyze the memory savings achieved by disaggregating the encoding and prefill stages.
+In this section, we analyze the memory savings achieved by disaggregating the encoding and prefill stages.
 The E workers can save memory as they do not require LLM weights or the KV cache.
 Analyzing only the weight size indicates memory reduction of approximately 95%, 96.2%, and 78.3% for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models.
-Similarly, for the P workers, memory saving of about 5%, 3.7%, and 21.6% are achieved, respectively.
+Similarly, for the P workers, memory savings of about 5%, 3.7%, and 21.6% are achieved, respectively.
 In practice, since KV cache is also not required at E workers, the memory saving can be even higher (93.3% saving, i.e., 15$\times$ lower, according to profiling).
-These reductions in memory usage enable the EPD system to support higher numbers of images per request and batch sizes (discussed in the following experiments) and larger KV cache sizes (discussed in Appendix D.2).
+These reductions in memory usage enable the EPD system to support higher numbers of images per request and batch sizes (discussed in the following experiments) and larger KV cache sizes (discussed in Appendix A.2).
 Model Image Reso. DistServe EPD
@@
 4032,3024 1 10
-Table 1: Comparison of maximum supported number of images/ request. Batch size is fixed to 1. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
+Table 2: Comparison of the maximum number of images supported per request for various image resolutions across different models. Higher values are better; best values in each row are italicized.
 EPD Supports a Higher Number of Images per Request: We compare the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B.
-The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and with 80% of the available memory allocated for KV cache.
-As shown in Tab.
-1, EPD handles more images per request than DistServe, for example, at 4032$\times$3024 resolution, 7$\times$ more for InternVL2-26B and 10$\times$ more for InternVL2-8B.
+The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and with 80% of the available memory allocated for the KV cache.
+As shown in Table 2, EPD handles more images per request than DistServe, for example, at 4032$\times$3024 resolution, 7$\times$ more for InternVL2-26B and 10$\times$ more for InternVL2-8B.
 For InternVL2-8B, the limit of 19 images per request is due to its maximum context length.
 Without this constraint, a larger number of images per request could be supported.
-EPD Supports Higher Batch Sizes: We compare maximum supported batch sizes for E and P stages across different settings of image resolution and models in Tab.
-2.
+EPD Supports Higher Batch Sizes: We compare maximum supported batch sizes for E and P stages across different settings of image resolution and models in Table 3.
+The number of images per request is fixed to 10, and the KV cache is allocated to utilize 80% of the available free memory.
 As seen, EPD significantly outperforms DistServe for both E and P batch sizes.
@@
 4032, 3024 13 OOM 6 1
-Table 2: Comparison of maximum supported E and P batch size. #images/ request is fixed to 10. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
+Table 3: Comparison of the maximum supported batch sizes for E and P stages across different models and image resolutions.
+Higher values are better; italicized values indicate the best in each row.
+OOM denotes cases where the model ran out of memory.
 4.4 Ablation Study
@@
 Effect of IRP: We analyze the effect of ablating IRP, presented in Section 3.2.2.
-The experimental settings are same as TTFT experiment in Sec.
-4.2.
-As shown in Tab.
-3, removing the IRP feature negatively affects TTFT across various images/ request.
+The experimental settings are the same as the TTFT experiment in Section 4.2.
+As shown in Table 4, removing the IRP feature negatively affects TTFT across various images/ request.
 Moreover, the degradation exacerbates as the number of images/ request increases.
@@
 w/o Opt. 0.56 (2.2x) 4.48 (2.1x) 0.025 (0.8x)
-Table 4: Effect of Offline Optimizer.
-When the configuration is set randomly, the goodput is 2.2× worse on average.
-$\downarrow$ indicates that lower is better and $\uparrow$ means that larger is better.
-Effect of Offline Optimizer: By default, the EPD system collects workload samples and finds the optimal config offline using the optimizer.
-To demonstrate its effect, we conduct an experiment without using the optimizer and set configs randomly.
-Specifically, we uniformly sample 10 configs at random and report the estimated expectation of the performance metric in the second row of Tab.
-4.
-For a fair comparison, when evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is the goodput of the EPD system.
-(See Appendix B.4 for more details) As shown in Tab.
-4, disabling the optimizer causes significant performance degradation, in terms of both goodput and TTFT.
-This demonstrates the necessity of the optimizer, especially when the EPD system has many tunable configs.
+Table 5: Ablating the offline optimizer reduces goodput by 2.2× on average when configurations are selected randomly. $\downarrow$ indicates lower is better; $\uparrow$ indicates higher is better.
+Effect of Offline Optimizer: By default, the EPD system collects workload samples and finds the optimal configuration offline using the optimizer.
+To demonstrate its effect, we conduct an experiment without using the optimizer and select configurations randomly.
+Specifically, we uniformly sample 10 configurations at random and report the expected value of the performance metric in the second row of Table 5.
+For a fair comparison, when evaluating TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which corresponds to the goodput of the EPD system (see Appendix E.4 for more details).
+As shown in Table 5, disabling the optimizer causes significant performance degradation in terms of both goodput and TTFT.
+This highlights the importance of the optimizer, especially when the EPD system has many tunable configurations.
 Effect of Dynamic Role Switching: To analyze the impact of dynamic role switching, we conduct a controlled experiment simulating a shift in workload characteristics.
 Specifically, we generate 100 requests following the same static workload configuration as described in Section 3.2.4.
-However, to introduce an artificial workload change, we set the first 10 requests to generate 50 output tokens, while the remaining 90 requests generate 500 output tokens.
+However, to introduce an artificial workload change, the first 10 requests generate 50 output tokens, while the remaining 90 requests generate 500 output tokens.
 The request arrival rate is fixed at 3 requests per second.
-The results are shown in Tab.
-5.
+The results are shown in Table 6.
 Latency (s) $\downarrow$ TTFT (s) $\downarrow$ TPOT (s) $\downarrow$
@@
 w/o Switch 61.10 (2.2x) 1.33 (0.9x) 0.12 (2.4x)
-Table 5: Effect of ablating dynamic role switching feature from EPD.
-Disabling the worker migration feature negatively affects the TPOT (2.4x worse), which results in significantly poor end-to-end latency (2.2x worse).
-Results are averaged over 100 requests with one 4K image per request.
-$\downarrow$ indicates that lower is better.
-Without dynamic worker migration, the system performs poorly because it remains locked in the initial configuration (5E1P2D, optimized offline for 50 tokens) and is unable to adapt to the increased decoding demand.
-In contrast, the EPD system with migration dynamically reconfigures itself (2E1P5D) to handle the new workload (500 output tokens) by shifting 3 E instances to D, resulting in approximately 2× better performance.
+Table 6: Ablating dynamic role-switching from EPD degrades TPOT by 2.4× and increases end-to-end latency by 2.2×.
+Results are averaged over 100 requests with one 4K image each.
+$\downarrow$ indicates lower is better.
+Without dynamic worker migration, the system performs poorly because it remains fixed in the initial configuration (5E1P2D, optimized offline for 50 tokens) and is unable to adapt to the increased decoding demand.
+In contrast, the EPD system with migration dynamically reconfigures itself (2E1P5D) to handle the new workload (500 output tokens) by shifting three E instances to D, resulting in approximately 2× better performance.
 4.5 Adaptation to Neural Processing Units (NPUs)
-This section presents the results of adapting the proposed EPD framework to Huawei Ascend NPUs. For more details on the implementation and experiments, see Sec. B.5.
-First, we compare the EPD-NPU against the vLLM and DistServe baselines in terms of SLO attainment using the InternVL2-8B LMM.
-We used the same settings as mentioned in Sec.
-4.1, except that we used a heavy encoding workload of 8 4032x3024 images per request.
-The optimal configuration for this workload was found to be 5E2P1D and the corresponding results are presented in Figure 8.
-As seen, EPD is the only setting that can achieve the SLO requirements, while the other baselines are completely decimated even for low request rates.
-Figure 8: SLO curve for InternVL2-8B with the following requirements (TTFT=8.5 and TPOT=0.12) on the controlled experimental setup with 8 images per request.
+We present the results of adapting the proposed EPD framework to Huawei Ascend NPUs. For more details on the implementation and experiments, see Appendix F.
+First, we compare EPD-NPU against the vLLM and DistServe baselines on the InternVL2-8B LMM.
+We evaluate SLO attainment, defined as TTFT $\leq$ 8.5s and TPOT $\leq$ 0.12s on eight 4K images per request.
+We used the same settings as described in Section 4.1, except that we employed a heavy encoding workload of eight $4032\times 3024$ images per request.
+The optimal configuration for this workload was found to be 5E2P1D, and the corresponding results are presented in Figure 9.
+As shown, EPD is the only configuration that achieves the SLO requirements, while the other baselines fail to meet the SLOs entirely, even at low request rates.
+Figure 9: SLO attainment ($\uparrow$) on NPUs under varying request rates for a synthetic workload using the InternVL2-8B model.
+EPD maintains positive SLO attainment under strict TTFT constraints, while baselines fail to meet the criteria.
 Next, we compare the TTFT of EPD-NPU against vLLM.
-We found that EPD-NPU demonstrated a 35.2% improvement over vLLM, which is significantly more than the 24.4% improvement for the GPU scenario.
-We attribute this additional $\sim$10to the proportionally heavier encode of LLMs on NPU compared to GPU.
-Please refer to Sec.
-D.3 for more details.
-This results in greater benefits of disaggregating the encode from prefill on NPU in comparison to GPU.
+EPD-NPU achieves a 35.2% improvement over vLLM, surpassing the 24.4% improvement observed in the GPU scenario.
+This additional $\sim$10improvement is attributed to the proportionally heavier encoding workload of LLMs on NPUs compared to GPUs.
+Please refer to Section F.1 for more details.
+As a result, disaggregating the encoding from the prefill phase yields greater benefits on NPUs than on GPUs.
 5 Conclusion
-In this paper, we present a novel approach to optimizing Large Multimodal Model (LMM) systems through the disaggregation of key processing stages.
-By separating the encoding, prefill, and decoding tasks into distinct stages, our system offers enhanced flexibility in resource allocation, enabling more efficient management of computational and memory resources.
-This disaggregation, combined with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage.
-Appendix A Optimizer Details
+In this paper, we proposed a novel approach for optimizing LMM systems via disaggregation of key processing stages.
+By separating the encoding, prefill, and decoding phases into distinct stages, our system provides greater flexibility in resource allocation, enabling more efficient management of computational and memory resources.
+This disaggregation, together with dynamic resource allocation, asynchronous token transfer, and advanced parallelization strategies, directly addresses several critical challenges in LMM deployment, including latency reduction, memory optimization, and efficient computational resource usage.
+We validated the effectiveness of our system through extensive experiments.
+Finally, we outline the system’s limitations and future directions in Appendix B and Appendix C, respectively.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Additional Experiments
+A.1 Adaptation to Audio modality
+In this section, we extend EPD to the audio modality to evaluate its effectiveness beyond vision tasks, highlighting the method’s generalizability.
+Specifically, we evaluate EPD in an online, audio setting using the ultravox-v0_3 model, based on LLaMA3.1-8B.
+To simulate an encode-intensive workload, each request is configured to include 24 audio files.
+Evaluation metrics include SLO attainment—defined as TTFT $\leq$ 2.0 s and TPOT $\leq$ 0.025 s—and goodput (requests per second, r/s).
+As shown in Table 7, EPD consistently outperforms vLLM and DistServe across all request rates, demonstrating superior efficiency and reliability.
+Combined with prior results in the video domain, these findings reinforce EPD’s robustness and generalizability across modalities.
+Rate (r/s) SLO Attainment Rates ($\uparrow$)
+vLLM DistServe EPD
+0.10 0.99 0.99 0.99
+0.25 1.00 0.94 0.99
+0.50 0.99 0.89 1.00
+1.00 0.91 0.72 0.96
+1.10 0.87 0.69 0.93
+1.15 0.87 0.68 0.93
+Goodput (r/s) $\uparrow$ 1.01 0.45 1.16
+Table 7: SLO attainment results ($\uparrow$) for online audio benchmarking with ultravox-v0_3 (24 audio files per request).
+All baselines use 4 GPUs: vLLM operates in data-parallel (DP) mode, DistServe uses a 3P1D configuration, and EPD adopts a 2E1P1D setup.
+EPD achieves consistently high SLO attainment and the highest goodput across all request rates.
+A.2 EPD Supports Larger KV Cache Sizes
+In this section, we compare the maximum KV cache size that can be allocated across various baselines.
+The batch size is fixed at 1, and the image resolution is set to 4032$\times$3024.
+As shown in Table 8, EPD supports significantly larger KV cache sizes than DistServe.
+For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for DistServe, representing a 2.2$\times$ improvement.
+Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an OOM error, indicating that it cannot process 20 images even with a KV cache size of 0.
+Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
+Model # Images/Req. DistServe EPD
+MiniCPM-V 2.6 5 86% 99%
+10 74% 97%
+20 49% 95%
+40 OOM 92%
+80 OOM OOCL
+InternVL2-8B 5 94% 95%
+10 89% 91 %
+20 OOCL OOCL
+InternVL2-26B 5 67% 89%
+10 36% 80%
+20 OOM 63%
+40 OOM OOCL
+Table 8: Comparison of maximum supported KV cache size (in terms of percentage of free memory) on prefill node for various #images/ request.
+Image resolution fixed to 4K.
+Higher (italicized) is better.
+A.3 Throughput Comparison for Offline and Heterogeneous scenario
+Figure 10: Left: Impact of varying the number of encoding workers in the EPD method.
+The notation xEyP denotes a configuration with x encoder and y prefill workers.
+The DistServe method uses a fixed 7P configuration, assigning 7 workers to handle both encoding and prefill steps.
+Middle: Effect of the number of images per request on end-to-end throughput.
+Right: Sensitivity to encoding and prefill batch sizes, where these two numbers are set equal.
+(a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
+Figure 11: SLO attainment ($\uparrow$) for end-to-end inference across multiple models and image counts per request.
+Subfigures (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively.
+The top and bottom rows show results for 6 and 8 images per request.
+EPD consistently outperforms all baselines, demonstrating robust performance as image count increases..
+In this section, we compare the throughput of the EPD and DistServe methods under both offline and heterogeneous settings.
+In the offline scenario, a batch of requests is submitted in advance, allowing the system to process them overnight with the goal of maximizing end-to-end (E2E) throughput.
+In the heterogeneous setting, we consider a cluster composed of GPUs with varying computational and memory capacities.
+Our motivation is as follows: in the DistServe system, both the encoding and prefill stages are executed on the same worker, requiring the corresponding memory to be allocated on a single GPU.
+However, GPU memory is limited, and this constraint becomes particularly problematic in heterogeneous environments that include low-end GPUs.
+In such cases, the total memory demand of the encoding and prefill stages may approach or exceed the available memory, significantly restricting the batch size or even rendering the DistServe method infeasible on these devices.
+Thus, we consider a controlled scenario in our experiment where the memory usage of the prefill stage in DistServe remains within the capacity of low-end GPUs, but only permits a minimal batch size of 1 for both encoding and prefill stages.
+We conduct this experiment using 8 A800 GPUs.
+For the EPD system, we use the default 5E2P1D configuration, which allocates 5, 2, and 1 GPUs to the encoding, prefill, and decoding stages, respectively.
+The corresponding maximum batch sizes are set to 8 for encoding, 8 for prefill, and 128 for decoding.
+In the DistServe system, we adopt the 7P1D setup, with maximum batch sizes of 1 and 128.
+Here, the encoding and prefill steps are conducted on the 7 P workers.
+The workload consists of 1,000 requests, each containing a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens.
+We perform an ablation study by tuning the system’s hyperparameters as described in Sec.
+3.2.3, and analyze their impact on throughput.
+The results are presented in Fig.
+10.
+The left plot demonstrates the importance of selecting an appropriate GPU configuration.
+The algorithm detailed in Sec.
+3.2.3 can automatically select the 5E2P configuration as the optimal setup for maximizing end-to-end (E2E) throughput.
+The middle plot shows that EPD achieves higher throughput when the number of images per request is small, indicating that its disaggregated architecture effectively mitigates encoder-side compute bottlenecks.
+The right plot highlights that EPD is relatively insensitive to encoding and decoding batch sizes.
+This enables users to either manually choose a large batch size or rely on the algorithm in Sec.
+3.2.3 to automatically determine an optimal configuration.
+A.4 SLO Attainment for Higher Number of Images
+In this section, we show additional results from the experiment in Section 4.1 pertaining to 6 and 8 images per request.
+The results are presented in Figure 11.
+As seen, the SLO attainment for EPD begins to decline as number of images rise, particularly at higher request rates.
+This decline reflects the increasing computational demand associated with processing multiple high-resolution images within a single request.
+Nevertheless, EPD continues to outperform all baselines, which struggle even at low request rates and fail to scale effectively under increased workloads.
+Appendix B Limitations
+While disaggregated serving offers significant benefits in meeting strict service-level objectives, it is not universally optimal.
+In scenarios where throughput is the primary concern and SLOs can be relaxed, traditional aggregated serving can outperform disaggregated approaches due to superior GPU utilization.
+This is largely attributed to the absence of inter-stage communication overhead and pipeline inefficiencies/ bubbles that can arise from resource imbalance in disaggregated setups.
+From a cost-efficiency perspective, disaggregated serving is most beneficial when latency SLOs are non-negotiable.
+In contrast, when maximizing throughput per dollar is the primary goal and latency constraints are minimal, aggregated configurations are often more cost-effective.
+This can be attributed to the reduced communication cost and orchestration overhead that result in higher GPU utilization.
+Specific to the EPD disaggregation strategy, the encoding and prefill stages are predominantly compute-bound.
+As a result, the primary performance gains from disaggregation stem from stage-specific finetuning of batching, scheduling, and parallelization strategies, especially IRP for parallelizing encoding load of a request across multiple GPUs.
+However, these gains are sensitive to pipeline inefficiencies, for instance pipeline bubbles can erode the latency benefits.
+Thus, careful tuning of resource allocation and responsive worker migration is required to ensure these stages are well-balanced for a given model and workload.
+Appendix C Future Work
+Although the encode and prefill stages have similar compute characteristics, their memory usage profiles differ substantially.
+This asymmetry presents a promising opportunity to explore heterogeneous hardware configurations.
+For instance, using high-memory high-compute GPUs for prefill and low-memory but high-compute units for encoding could further enhance efficiency and cost-effectiveness in disaggregated setups.
+Another area of exploration is reducing the overhead of visual token migration, which can be a bottleneck in multimodal disaggregated pipelines.
+Compression techniques, particularly those leveraging the inherent redundancy in visual tokens as identified in token pruning literature (Alvar et al., 2025), may significantly reduce transfer latency and cost.
+Lastly, disaggregated serving introduces new possibilities in privacy-aware inference (Liu et al., 2020; Huang et al., 2020; Singh et al., 2022), especially in edge-cloud environments.
+For example, encoding could be performed on edge devices, keeping raw, privacy-sensitive images local.
+Only the encoded representations would be transmitted to the cloud for prefill and decode stages, minimizing privacy risks without compromising model capability.
+This hybrid architecture could form the foundation for privacy-preserving multimodal systems.
+Appendix D Optimizer Details
 We recapitulate the configuration optimization problem:
@@
 Second, how should the order of requests be determined within a worker queue?
-To solve these decisions, we adapt strategies from the DistServe framework [21].
+To solve these decisions, we adapt strategies from the DistServe framework (Zhong et al., 2024).
 In the encoding stage, when a request arrives, it is assigned and pushed to an instance queue.
@@
 For simplicity, we constrain that all instances within the same stage share the same scheduling strategy.
-Appendix B Implementation Details
+(a) GPU                                   (b) NPU
+Figure 12: Breakdown of latency for encode and prefill stages using the InternVL2-8B model across varying numbers of images per request.
+Subfigures (a) and (b) show results on GPU and NPU, respectively.
+Light green denotes encode latency and light blue indicates prefill latency.
+NPUs demonstrate distinct latency characteristics compared to GPUs as input size increases.
+Appendix E Implementation Details
 EPD is a fully capable distributed serving system for LMMs, comprising several key components: a load estimation module, a resource allocation module, a RESTful API frontend, and a multimodal-aware orchestration layer.
@@
 For efficient data movement, the system employs NCCL for inter-node GPU communication and asynchronous CudaMemcpy for intra-node transfers, ensuring smooth operations without disrupting GPU computations across the disaggregated EPD framework.
-B.1 Hyper-parameters in EPD system
+E.1 Hyper-parameters in EPD system
 We conducted our experiments using a cluster of 8 NVIDIA A100 GPUs (82GB).
@@
 Further, to optimize TTFT and TPOT in this latency-sensitive setting, we disabled batching by setting batch sizes to 1 for the encoding, prefill, and decoding stages.
-B.2 LMM Details
+E.2 LMM Details
 MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in 8 billion parameters.
-It excels in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [19].
+It excels in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains (Yao et al., 2024).
 InternVL2-8B features an InternViT-300M-448px vision encoder with 300 million parameters, paired with an internlm2_5-7b-chat language model comprising 7.7 billion parameters, totaling 8 billion parameters.
 Similarly, InternVL2-26B combines an InternViT-6B-448px-V1-5 vision encoder with 6 billion parameters and an internlm2-chat-20b language model containing 20 billion parameters, resulting in 26 billion parameters.
-B.3 SLO Criteria
-Tab.
-6 outlines the SLO criteria for TTFT and TPOT across various models and different numbers of images per request (#I/R).
+E.3 SLO Criteria
+Table 9 outlines the SLO criteria for TTFT and TPOT across various models and different numbers of images per request (#I/R).
 These criteria are empirically derived based on the characteristics of the underlying models, such as the computational complexity of the MME and LLM.
@@
 In contrast, TPOT requires only minor adjustments since it is not directly impacted by changes in #I/R.
-MiniCPMv 2.6 8B InternVL 8B InternVL 26B
-#I/R TTFT TPOT TTFT TPOT TTFT TPOT
+#I/R MiniCPM-V 2.6 InternVL 8B InternVL 26B
+TTFT TPOT TTFT TPOT TTFT TPOT
 2 1.40 0.04 1.20 0.05 3.50 0.07
@@
 8 5.10 0.06 5.00 0.18 15.00 0.15
-Table 6: Requirements for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
-B.4 Details for the Ablation Experiment of Offline Optimizer
+Table 9: TTFT and TPOT values (in seconds) used as SLO thresholds for different models and image counts per request (#I/R).
+Respective values are shown for MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models.
+E.4 Details for the Ablation Experiment of Offline Optimizer
 In this experiment, 100 User requests arrive in real-time with each request containing 6 images, and they all request for the MiniCPMv model.
@@
 For a fair comparison, When evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is same as the goodput of the EPD system.
-B.5 NPU Implementation Details
-There are intrinsic differences between NPU and GPU architectures and processing characteristics.
-GPUs are general-purpose compute resources whereas NPUs have been specifically designed for neural networks.
-NPUs are optimized for FP16 calculations whereas GPUs are optimized for FP32 calculations.
-As stated above, CANN operators are used on Ascend NPU hardware in place of CUDA operators.
-These differences between CUDA and corresponding CANN operators, having been optimized for the hardware they are built upon, result in different latencies and trade-offs in inference as well as communication between NPU/GPU chips.
-This results in different encode/prefill/decode latencies which may be present different benefits from EPD and IRP.
-The EPD-NPU system was implemented in a similar manner to our EPD-GPU system as described in Appendix B.
+Appendix F Extending EPD to NPU
+We extend our EPD framework to a cluster of Neural Processing Units (NPUs) for two purposes: (1) to investigate the impact of disaggregating (E) from (P) on NPUs which tend to have higher encode-to-decode latency ratio and (2) to demonstrate the generalizability of EPD.
+F.1 NPUs having higher encode-to-prefill latency ratios
+When investigating architectures and potential benefits of an EPD-NPU framework, we first profiled the LMMs tested in this paper for their encoding and prefilling latencies to understand the impact of the characteristics described above.
+We calculated the average encode and prefill stage latencies of 10 requests.
+Across the different LMMs, NPUs spent more time on encode than prefill across these models when compared to GPUs.
+Figure 12 shows the latency breakdown between encode and prefill for InternVL2-8B across a range of #I/R.
+Across the models and different #I/R, we found a $\sim$10-20larger encode-to-prefill latency ratios in NPU than in GPU.
+This trend was consistent across different large multimodal models (LMMs), indicating that NPUs tend to spend a greater proportion of time in the encoding phase compared to GPUs.
+Since the encode in NPU is a larger portion of the TTFT when compared to GPU, we hypothesized that we would see more benefit from EPD and IRP towards NPU. Thus, we were motivated to adapt EPD for NPU.
+F.2 Implementation Details for NPU
+The EPD-NPU system was implemented in a similar manner to our EPD-GPU system as described in Appendix E.
 All EPD-NPU experiments utilize IRP.
 The EPD-NPU distributed serving system was developed and deployed on a cluster of 910B3 NPUs, each with 64GB of high bandwidth memory.
-Some primary differences between the two implementations include:
+There are two major differences between the NPU and GPU implementations:
   • Ascend-vLLM: EPD-NPU leverages Ascend-vLLM rather than vLLM which is an adaptation of vLLM on Ascend NPU hardware that utilizes the Compute Architecture for Neural Networks (CANN) software stack rather than CUDA.
@@
 This container-based deployment allows our framework to be easily scalable and portable to the cloud, providing intriguing future works regarding efficient utilization and scheduling of resources for users.
-Figure 9: Left: The impact of tuning the number of encoding workers in the EPD method.
-The annotation ‘xEyP’ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P’ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests.
-Middle: The impact of the number of images per request in the workload.
-Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
-These differences in characteristics provide an opportunity to investigate the generalizability of our EPD framework and further understand the optimal setting/environment for disaggregating encode and prefill.
+F.3 Experimental Settings for NPU
 For EPD-NPU, We use eight 910B3 NPUs and evaluate the system by running the controlled experimental setup (as described in the experiments section of the main paper) in an online setting.
-Specifically, the number of input tokens was 22, length of output tokens was 10 with early stopping was disabled.
-The SLO requirements were also selected in a similar manner as the GPU experiments.
-The requirements for InternVL2-8B with 8 images per request were 8.5 and 0.12 for TTFT and TPOT, respectively.
-Each server had 192 CPUs and 1.5TB of RAM.
-The CANN version was 7.6.
-The maximum model length was set to 27000, KV Cache utilization was set to 83% And the maximum number of multimedia data of 8 was imposed.
-EPD-NPU also used a version of Ascend-vLLM (version 0.6.3.post1) built upon a newer version of vLLM than the one used in EPD-GPU.
-Unless otherwise stated, the hyper-parameters are identical to those described in B.1.
-Appendix C Throughput in Offline Settings
-In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the E and P stages.
-Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput.
-This is an additional performance metric complementing the main body of the paper, measured in terms of requests completed per second and the system’s ability to handle high request volumes in offline settings.
-In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
-This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory.
-In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory.
-In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory.
-In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
-(a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
-Figure 10: Columns (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively.
-The top and bottom rows show results for 6 and 8 images per request, respectively.
-As expected, EPD maintains a significant performance advantage over all baselines, demonstrating robustness of the system as the image count increases.
-We use 8 A800 GPUs for this experiment.
-For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively.
-In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively.
-The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens.
-We tune the system’s hyperparameters as outlined in Sec.
-3.2.3 and evaluate their impact on throughput.
-The results are presented in Fig.
-9.
-The left plot illustrates the importance of selecting an appropriate GPU configuration.
-As described in Sec.
-3.2.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput.
-The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck.
-The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method.
-This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec.
-3.2.3 to automatically select an optimal batch size.
-Appendix D Additional Experiments
-D.1 SLO Attainment for End-to-End Generation for 6 and 8 images
-In this section, we show additional results from the experiment in Sec.
-4.1 pertaining to 6 and 8 images per request.
-As seen, the SLO attainment for EPD begins to decline as number of images rise, particularly at higher request rates.
-This decline reflects the increasing computational demand associated with processing multiple high-resolution images within a single request.
-Nevertheless, EPD continues to outperform all baselines, which struggle even at low request rates and fail to scale effectively under increased workloads.
-Model # Images/Req. DistServe EPD
-MiniCPM-V 2.6 5 86% 99%
-10 74% 97%
-20 49% 95%
-40 OOM 92%
-80 OOM OOCL
-InternVL2-8B 5 94% 95%
-10 89% 91 %
-20 OOCL OOCL
-InternVL2-26B 5 67% 89%
-10 36% 80%
-20 OOM 63%
-40 OOM OOCL
-Table 7: Comparison of maximum supported KV cache size (in terms of percentage of free memory) on prefill node for various #images/ request.
-Image resolution fixed to 4K.
-Higher (italicized) is better.
-(a) GPU                        (b) NPU
-Figure 11: Breaking down the latency of encode and prefill for InternVL2-8B across different number of images per requests.
-We compare the difference between GPU (left) and NPU (right) where light green is encode and light blue is prefill.
-D.2 EPD Supports Larger KV Cache Sizes:
-We compare the maximum KV cache size that can be allocated across systems.
-The batch size is fixed at 1, and the image resolution is set to 4032$\times$3024.
-As shown in Tab.
-7, EPD supports significantly larger KV cache sizes than DistServe.
-For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for DistServe, representing a 2.2$\times$ improvement.
-Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an OOM error, indicating that it cannot process 20 images even with a KV cache size of 0.
-Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
-D.3 Heavier Encode in NPUs
-We compare the encode-to-prefill stage latencies between NPUs and GPUs.
-Figure  11 show the results for the InternVL2-8B model, across various I/R values.
-As seen, we notice higher ($\approx 10-20\%$) encode-to-prefill latency ratios for NPUs in comparison to GPUs, indicating the need for a different optimized configuration for NPU.
-This trend was consistent across different large multimodal models (LMMs), indicating that NPUs tend to spend a greater proportion of time in the encoding phase compared to GPUs.
+The number of input tokens was 22, length of output tokens was 10 with the end-of-sequence token ignored.
+The SLO requirements (8.5 and 0.12 for TTFT and TPOT, respectively) were also selected in a similar manner as the GPU experiments.
+The maximum model length was set to 27000, and KV Cache utilization was set to 83%.
+EPD-NPU also used a version of Ascend-vLLM (version 0.6.3.post1) with CANN version 7.6.
+Unless otherwise stated, the hyper-parameters are identical to those described in E.1.
+We evaluate EPD-NPU using InternVL2-8B LMM against two baselines: (1) Ascend-vLLM and (2) Ascend-vLLM PD disaggregation for the heavy workload of 8 images (4032x3024 resolution) per request.